### v2

# Capacity Bounds for Broadcast Channels with Bidirectional Conferencing Decoders

Abstract
The two-user broadcast channel (BC) with receivers connected by bidirectional cooperation links of finite capacities, known as conferencing decoders, is considered. A novel capacity region outer bound is established based on multiple applications of the Csiszár-Körner identity. Achievable rate regions are derived by using Marton’s coding as the transmission scheme, together with different combinations of decode-and-forward and quantize-bin-and-forward strategies at the receivers. It is shown that the outer bound coincides with the achievable rate region for a new class of semi-deterministic BCs with degraded message sets; for this class of channels, one-round cooperation is sufficient to achieve the capacity. Capacity result is also derived for a class of more capable semi-deterministic BCs with both common and private messages and one-sided conferencing. For the Gaussian BC with conferencing decoders, if the noises at the decoders are perfectly correlated (i.e., the correlation is either 1 or -1), the new outer bound yields exact capacity region for two cases: i) BC with degraded message sets; ii) BC with one-sided conferencing from the weaker receiver to the stronger receiver. When the noises have arbitrary correlation, the outer bound is shown to be within half a bit from the capacity region for these same two cases. Finally, for the general Gaussian BC, a one-sided cooperation scheme based on decode-and-forward from the stronger receiver to the weaker receiver is shown to achieve the capacity region to within $\frac{1}{2}\log(\frac{2}{1-|\lambda|})$ bits, where $\lambda$ is the noise correlation. An interesting implication of these results is that for a Gaussian BC with perfectly negatively correlated noises and conferencing decoders with finite cooperation link capacities, it is possible to achieve a strictly positive rate using only infinitesimal amount of transmit power.
Index Terms:
Broadcast channel, bidirectional cooperation, conferencing decoders, capacity region, deterministic channel, Gaussian broadcast channel
I Introduction
In many communication networks, it is sometimes feasible for the receivers at distinct locations to exchange messages and to cooperate. Receiver cooperation is known to be able to improve the performance of the direct communication from the transmitter. This paper considers a channel model in which cooperation takes place via dedicated digital links of finite capacities—termed conferencing links. We investigate the impact of receiver cooperation on the capacity region of a two-user broadcast channel (BC) with conferencing links, with both common and private messages.
I-A Motivation
We are particularly interested in the effect of receiver noise correlation on the capacity of BC with conferencing decoders. To motivate the discussion, consider a Gaussian channel model as shown in Fig. 1. As is well known, the capacity of a single-user Gaussian channel is $C=\frac{1}{2}\log\left(1+\frac{P}{N}\right)$ bits per channel use, where $P$ is the average transmit power constraint and $N$ is the receiver noise power. Thus, at a fixed positive noise power $N$, if the transmit power $P\to 0$, then $C\to 0$.
Now consider the scenario of transmitting a common message through a two-user Gaussian broadcast channel with conferencing decoders. If the noises at the two receivers have the same power and they are uncorrelated, then full receiver cooperation can potentially increase the signal-to-noise ratio (SNR) by 3dB, resulting in up to 1 bit increase in the common message capacity of the BC.
The situation is completely different if the receiver noises are correlated. Specifically, if the receiver noises are perfectly positively correlated, then cooperation would gain nothing, because $Y_{1}=Y_{2}$ in this case; as each receiver already knows what the other receiver knows, exchanging information does not help. However, if the receiver noises are perfectly negatively correlated, then cooperation can help significantly. Consider the case where $Y_{1}=X+Z$ and $Y_{2}=X-Z$, and the conferencing links have infinite capacities $C_{12}=C_{21}=\infty$. Then, the BC with conferencing decoders would have infinite capacity at any positive transmit power $P$, because having access to both $Y_{1}$ and $Y_{2}$ allows the noises to be cancelled completely.
What happens if the noises are perfectly negatively correlated, but the conferencing links have finite capacity, e.g., $C_{12}=C_{21}=1$ bit per channel use? The analysis becomes more delicate and this is one of the main topics of this paper. It turns out that an interesting phenomenon emerges—it is possible to achieve a strictly positive rate (i.e., 1 bit in the example above), even with an infinitesimally small transmit power. In contrast to the single-user Gaussian channel where $P\to 0$ implies $C\to 0$, for this Gaussian BC with perfectly negatively correlated noises and conferencing decoders of fixed cooperative link capacities, we can have $P\to 0$, while $C\to 1$. In other words, the energy efficiency for transmitting one bit of information can be infinite!
Figure 1: The capacity of a Gaussian channel (left) goes to zero if the transmit power $P$ goes to zero at a fixed noise variance $N$. In a Gaussian BC with perfectly negatively correlated noises at the receiver (right), conferencing links of fixed capacities $C_{12}$ and $C_{21}$ between the receivers can allow a strictly positive rate to be achieved even at an infinitesimally small transmit power.
The above example motivates us to study the BC with conferencing decoders with a specific focus on the Gaussian channel with correlated noises. We remark that correlated noises often occur in practical communication scenarios when the noises at the two receivers are affected by a strong common interference source.
I-B Prior Literature
Cooperation via conferencing links has been studied for many different communication networks in the information theory literature [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Further, the papers [19, 20, 21, 22, 23, 24, 25, 26] specifically consider the two-user BC with conferencing decoders. In [19], the authors develop communication strategies for the interactive decoding of a common message broadcast to cooperative receivers. In [20], the capacity region of physically degraded channel is derived and also an achievable rate region is given for the general case. In [21], an achievable rate region is presented based on coding strategies for the partially cooperative relay broadcast channels and also a converse result is proved. In [22], the problems of communication over physically degraded, state-dependent BCs with one-sided conferencing decoders are investigated. In [23], the capacity region of the semi-deterministic BC with one-sided decoder cooperation is derived and its duality with a source coding problem is addressed. The authors in [24] consider the BC with one-sided cooperating receivers under the strong secrecy constraints and present capacity results for semi-deterministic and physically degraded cases. In [25], the BC with (one-sided) unreliable cooperating decoders is studied. In a recent work [26], the BC with degraded message sets and one-sided cooperation link that may be absent is considered and its capacity region is given.
I-C Main Contributions
The existing capacity results for the BC with conferencing decoders are all for the case of one-sided cooperation, i.e., only one of the decoders is connected to the other by a cooperating link. This is due to the lack of useful outer bounds (beyond the cut-set bound) for the two-sided cooperation case. In this paper, we first establish a novel outer bound on the capacity region of the two-user BC with bidirectional conferencing decoders. The new outer bound, which is derived using multiple applications of the Csiszár-Körner identity [27] [28, Lemma 7], is strictly tighter than the previous ones including that of [20, Proposition 1], which is essentially the cut-set bound.
We then propose achievability strategies for the BC with conferencing decoders. In [20, Theorem 2] an achievable region is derived for the BC with bidirectional cooperation by applying Marton’s coding at the transmitter and the compress-and-forward cooperative scheme at both receivers. Another achievable region is given in [21] based on coding strategies for the partially cooperative relay broadcast channels. A third achievable region is given in [23, Appendix B] for the BC with one-sided cooperation between receivers, which is derived by applying Marton’s coding at the transmitter and decode-and-forward as the cooperative protocol. All of these achievable schemes are however insufficient to either derive new capacity results for the discrete memoryless BCs, or approximate capacity results for the Gaussian BC.
This paper presents two achievability schemes for the two-user BC with both common and private messages and bidirectional conferencing decoders. In the first scheme, we apply Marton’s coding as the transmission scheme, and quantize-bin-and-forward at one receiver first and then a combination of decode-and-forward and quantize-bin-and-forward at the other receiver as cooperative strategy. In the second scheme, we apply a combination of decode-and-forward and quantize-bin-and-forward at one receiver first and then quantize-bin-and-forward at the other receiver as cooperation scheme. For each achievability scheme, at the receiver that applies a combination of decode-and-forward and quantize-bin-and-forward, the optimal proportion of the capacity of the cooperative link that should be devoted to each strategy is determined.
We prove that the novel outer bound coincides with the first achievable rate region for a new class of semi-deterministic BCs with degraded message sets. This capacity result is important from two viewpoints. First, it is the first capacity result for the two-user BC with two-sided conferencing decoders (all previously known capacity results are regarding the channel with one-sided cooperation). Second, it is among the rare cases in network information theory for which quantize-bin-and-forward is optimal. These results also demonstrate that a one-round cooperation protocol is sufficient to achieve capacity and multi-round strategies similar to those devised in [19] are not needed. Moreover, we derive a capacity result for a new class of more capable semi-deterministic BCs with both common and private messages and one-sided cooperation.
Furthermore, we evaluate the derived outer bound for the Gaussian channel with correlated noises. Using this bound, we prove several interesting results. For the channel with perfectly correlated noises (when noise correlation is either 1 or -1), the new outer bound yields the exact capacity region for two cases: i) BCs with degraded message sets; ii) BCs with one-sided conferencing between decoders. For these two cases, we also show that the outer bound is within half a bit from the capacity region over arbitrary noises correlation. Lastly, we prove that for the Gaussian BC, a one-sided cooperative scheme from the stronger receiver to the weaker receiver based on the decode-and-forward technique is already sufficient to achieve the capacity region to within $\frac{1}{2}\log(\frac{2}{1-\left|\lambda\right|})$ bits, where $\lambda$ is the correlation coefficient of the noises at the receivers of the BC. Therefore, such a strategy is approximately optimal when the noise correlation is small.
I-D Organization of Paper and Notations
The rest of the paper is organized as follows. We begin by defining the channel model in Section II. The main converse and achievability results for the discrete memoryless BC are presented in Section III. Section IV treats the Gaussian BC with conferencing decoders. Section V concludes the paper. The appendices contain most of the proofs.
In this paper, we use the following notations. A random variable is given by upper case letter (e.g. $X$) and its realization is shown by lower case letter (e.g. $x$). We use $X^{n}$ to denote a sequence of random variables $(X_{1},\cdots,X_{n})$ and use the notation $X_{t}^{n}=(X_{t},X_{t+1},\cdots,X_{n})$. Further, for integers $m$ and $n$, we use $[m:n]$ to denote the index set $\{m,\cdots,n\}$.
The probability density function (PDF) or the probability mass function (PMF) of a random variable $X$ is denoted by $P_{X}(x)$ and the conditional PDF/PMF of $X$ given $Y$ is denoted by $P_{X|Y}(x|y)$, where the subscripts are omitted occasionally for brevity. The operator $\{a\}^{+}$ is defined as: $\{a\}^{+}=\max\{0,a\}$. The set of non-negative real numbers is given by $\mathbb{R}_{+}$. For $\epsilon>0$, the set of all jointly $\epsilon$-letter typical $n$-sequences $x^{n}$ with respect to the PDF/PMF $P_{X}(x)$ is denoted by $\mathcal{T}^{n}_{\epsilon}(P_{X})$; (see [29] for definition). We use the notation $X=\varnothing$ to denote that a random variable $X$ is set as a constant. Finally, we use the shorthand $\psi(x)=\frac{1}{2}\log(1+x)$.
Figure 2: Two-user BC with conferencing decoders
II Channel Model
The two-user BC with conferencing decoders is a communication scenario in which a transmitter sends a common message and two private messages to two receivers, and the two receivers are able to exchange information via communication links of finite capacities (called conferencing links). Fig. 2 illustrates the channel model. The channel is given by $(\mathcal{X},\mathcal{Y}_{1},\mathcal{Y}_{2},P(y_{1},y_{2}|x),C_{12},C_{21})$ where $\mathcal{X}$ denotes the input alphabet, $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$ denote the output alphabets, $P(y_{1},y_{2}|x)$ is the channel transition probability, and $C_{12}$ and $C_{21}$ are capacities of the conferencing links per channel use.
For the BC with conferencing decoders, a length-$n$ code with $\Gamma$ conferencing rounds is described as follows. The transmitter encodes independent messages $M_{0}$, $M_{1}$, and $M_{2}$, which are uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively, into a codeword and sends over the channel according to the following:
$\Delta:[1:2^{nR_{0}}]\times[1:2^{nR_{1}}]\times[1:2^{nR_{2}}]\longmapsto\mathcal{X}^{n}$
$X^{n}=\Delta(M_{0},M_{1},M_{2}).$
The receiver $Y_{i},\;i=1,2$, receives a sequence $Y_{i}^{n}\in\mathcal{Y}_{i}^{n}$. The code consists of two sets of conferencing functions $\{\Xi_{12,\gamma}\}_{\gamma=1}^{\Gamma}$ and $\{\Xi_{21,\gamma}\}_{\gamma=1}^{\Gamma}$ with the corresponding output alphabets $\{\mathcal{J}_{12,\gamma}\}_{\gamma=1}^{\Gamma}$ and $\{\mathcal{J}_{21,\gamma}\}_{\gamma=1}^{\Gamma}$, respectively, which are described as follows:
$\Xi_{12,\gamma}$ $:\mathcal{Y}_{1}^{n}\times\mathcal{J}_{21,1}\times...\times\mathcal{J}_{21,\gamma-1}\longmapsto\mathcal{J}_{12,\gamma}$
$J_{12,\gamma}=\Xi_{12,\gamma}(Y_{1}^{n},J_{21}^{\gamma-1})$
$\Xi_{21,\gamma}$ $:\mathcal{Y}_{2}^{n}\times\mathcal{J}_{12,1}\times...\times\mathcal{J}_{12,\gamma-1}\longmapsto\mathcal{J}_{21,\gamma}$
$J_{21,\gamma}=\Xi_{21,\gamma}(Y_{2}^{n},J_{12}^{\gamma-1}).$
A sequence of conferencing rounds is said to be $(C_{12},C_{21})$-permissible if
$\sum_{\gamma=1}^{\Gamma}\log|\mathcal{J}_{12,\gamma}|\leq nC_{12},\qquad\sum_{\gamma=1}^{\Gamma}\log|\mathcal{J}_{21,\gamma}|\leq nC_{21}.$
Before decoding, the receivers exchange information by holding a $(C_{12},C_{21})$-permissible conference. The first receiver obtains the sequence $J_{21}^{\Gamma}=(J_{21,1},J_{21,2},...,J_{21,\Gamma})$ and the second one obtains the sequence $J_{12}^{\Gamma}=(J_{12,1},J_{12,2},...,J_{12,\Gamma})$. The receivers then decode their respective messages based on the following decoding functions:
$\nabla_{1}$ $:\mathcal{Y}_{1}^{n}\times\mathcal{J}_{21}^{\Gamma}\longmapsto[1:2^{nR_{0}}]\times[1:2^{nR_{1}}]$
$(\hat{M}_{0},\hat{M}_{1})=\nabla_{1}(Y_{1}^{n}\times J_{21}^{\Gamma})$
$\nabla_{2}$ $:\mathcal{Y}_{2}^{n}\times\mathcal{J}_{12}^{\Gamma}\longmapsto[1:2^{nR_{0}}]\times[1:2^{nR_{2}}]$
$(\hat{M}_{0},\hat{M}_{2})=\nabla_{2}(Y_{2}^{n}\times J_{12}^{\Gamma}).$
The capacity region for the channel is defined as usual [29]. Here, we omit the details for brevity.
III Discrete Memoryless BC with Conferencing Decoders
III-A Converse
Theorem 1.
Consider the two-user BC with conferencing decoders shown in Fig. 2. Let $\mathcal{R}_{o}$ denote the set of all rate triples $(R_{0},R_{1},R_{2})$ such that
$R_{0}+R_{1}$ $\leq I(U;Y_{1})+C_{21}$ (1)
$R_{1}$ $\leq I(X;Y_{1}|Y_{2},V)+I(X;Y_{2})$ (2)
$R_{1}$ $\leq I(X;Y_{2}|Y_{1},V)+I(X;Y_{1})$ (3)
$R_{0}+R_{2}$ $\leq I(V;Y_{2})+C_{12}$ (4)
$R_{2}$ $\leq I(X;Y_{2}|Y_{1},U)+I(X;Y_{1})$ (5)
$R_{2}$ $\leq I(X;Y_{1}|Y_{2},U)+I(X;Y_{2})$ (6)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{1}|V)+I(V;Y_{2})+C_{12}+C_{21}$ (7)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{2}|U)+I(U;Y_{1})+C_{12}+C_{21}$ (8)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{1}|Y_{2},V)+I(X;Y_{2})+C_{12}$ (9)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{2}|Y_{1},U)+I(X;Y_{1})+C_{21}$ (10)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{1},Y_{2})$ (11)
for some joint PDFs $P(u,v,x)$, where $U,V\rightarrow X\rightarrow Y_{1},Y_{2}$ forms a Markov chain and $|\mathcal{U}|\leq|\mathcal{X}|+2$ and $|\mathcal{V}|\leq|\mathcal{X}|+2$. The set $\mathcal{R}_{o}$ constitutes an outer bound on the capacity region.
Proof:
The proof is given in Appendix A. ∎
This novel outer bound on the capacity region of the channel is derived through multiple applications of the Csiszár-Körner identity [27] [28, Lemma 7]. The Csiszár-Körner identity is used to establish the constraints (7), (8), (9), and (10) based on the appropriate definitions of auxiliary random variables $U$ and $V$.
Specifically, the definitions of $U$ and $V$ are carefully chosen to ensure a compact representation of the outer bound using only two auxiliary random variables. Moreover, the multi-letter mutual information expressions are carefully manipulated in order to establish novel structures in the constraints (2), (3), (5), (6), (9) and (10) in terms of these auxiliary variables. Note that the constraint (11) is due to the cut-set bound.
As we demonstrate later, the novel structure of the outer bound is crucial for deriving new capacity results for specific channels (and also approximate capacity results for specific Gaussian channels).
The outer bound $\mathcal{R}_{o}$ is tighter than that of [20, Proposition 1], which is essentially the cut-set bound.
III-B Achievability
We now present novel achievability schemes for the two-user BC with both common and private messages and bidirectional conferencing decoders. The achievability schemes consist of judicious combination of Marton’s coding together with decode-and-forward and quantize-bin-and-forward cooperation strategies. The achievable rate regions are given in the following results.
Proposition 1.
Consider the two-user BC with conferencing decoders shown in Fig. 2. Let $\mathcal{R}_{i}^{(1)}$ denote the set of all rate triples $(R_{0},R_{1},R_{2})\in\mathbb{R}_{+}^{3}$ such that
$R_{0}+R_{1}$ $\leq\min\Big{\{}I(U,W;Y_{1})+\zeta_{2},I(U,W;Y_{1},\hat{Y}_{2})\Big{\}}$ (12)
$R_{0}+R_{2}$ $\leq\min\Big{\{}I(V,W;Y_{2})+\zeta_{1}+\bar{\alpha}_{1}C_{12},$
$\qquad I(V,W;\hat{Y}_{1},Y_{2})+\bar{\alpha}_{1}C_{12}\Big{\}}$ (13)
$R_{0}+R_{1}+R_{2}$ $\leq\min\Big{\{}I(U;Y_{1}|W)+\zeta_{2},I(U;Y_{1},\hat{Y}_{2}|W)\Big{\}}$
$\quad+\min\Big{\{}I(V,W;Y_{2})+\zeta_{1}+\bar{\alpha}_{1}C_{12},$
$\qquad\qquad\qquad I(V,W;\hat{Y}_{1},Y_{2})+\bar{\alpha}_{1}C_{12}\Big{\}}$
$\quad-I(U;V|W)$ (14)
$R_{0}+R_{1}+R_{2}$ $\leq\min\Big{\{}I(U,W;Y_{1})+\zeta_{2},I(U,W;Y_{1},\hat{Y}_{2})\Big{\}}$
$\quad+\min\Big{\{}I(V;Y_{2}|W)+\zeta_{1},$
$\qquad I(V;\hat{Y}_{1},Y_{2}|W)\Big{\}}-I(U;V|W)$ (15)
$2R_{0}+R_{1}+R_{2}$ $\leq\min\left\{I(U,W;Y_{1})+\zeta_{2},I(U,W;Y_{1},\hat{Y}_{2})\right\}$
$\quad+\min\Big{\{}I(V,W;Y_{2})+\zeta_{1}+\bar{\alpha}_{1}C_{12},$
$\qquad\qquad\qquad I(V,W;\hat{Y}_{1},Y_{2})+\bar{\alpha}_{1}C_{12}\Big{\}}$
$\quad-I(U;V|W)$ (16)
where
$\zeta_{1}$ =$=\{\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (17)
$\zeta_{2}$ =$=\{C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (18)
for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$ and $\alpha_{1}\in[0,1]$. The convex closure of the set $\mathcal{R}_{i}^{(1)}$ is achievable.
Proof:
The cooperation strategy to derive the rate region $\mathcal{R}_{i}^{(1)}$ involves using Marton’s coding scheme for the BC together with quantize-bin-and-forward at the receiver $Y_{2}$ first and then a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{1}$. At the receiver $Y_{1}$, the capacity of the conferencing link $C_{12}$ is divided into two parts. A fraction $\bar{\alpha}_{1}C_{12}$ is devoted for decode-and-forward and the remaining $\alpha_{1}C_{12}$ for quantize-bin-and-forward scheme. A detailed proof of Proposition 1 is given in Appendix B. ∎
Remark 1.
In the characterization of the rate region $\mathcal{R}_{i}^{(1)}$ given in Proposition 1, one can replace $\zeta_{1}$ and $\zeta_{2}$ with $\tilde{\zeta}_{1}$ and $\tilde{\zeta_{2}}$, respectively, as given below:
$\tilde{\zeta}_{1}$ $=\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})$ (19)
$\tilde{\zeta_{2}}$ $=C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}).$ (20)
To see why this is the case, it is sufficient to note that in the case of $\alpha_{1}C_{12}<I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})$, we can set $\hat{Y}_{1}=\varnothing$ and obtain the same rate region $\mathcal{R}_{i}^{(1)}$. Similarly, in the case of $C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, we can set $\hat{Y}_{2}=\varnothing$ and obtain the same rate region.
$R_{0}+R_{1}$ $\leq\min\left\{I(U,W;Y_{1})+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}),I(U,W;Y_{1},\hat{Y}_{2})\right\}$ (21)
$R_{0}+R_{2}$ $\leq I(V,W;Y_{2})+C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})$ (22)
$R_{0}+R_{1}+R_{2}$ $\leq\min\left\{I(U;Y_{1}|W)+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}),I(U;Y_{1},\hat{Y}_{2}|W)\right\}$
$\qquad+I(V,W;Y_{2})+C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})-I(U;V|W)$ (23)
$R_{0}+R_{1}+R_{2}$ $\leq\min\left\{I(U,W;Y_{1})+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}),I(U,W;Y_{1},\hat{Y}_{2})\right\}$
$\qquad+\min\left\{\begin{array}[]{l}{I(V;Y_{2}|W)+C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2}),}\\ {I(V;\hat{Y}_{1},Y_{2}|W)}\end{array}\right\}-I(U;V|W)$ (26)
$2R_{0}+R_{1}+R_{2}$ $\leq\min\left\{I(U,W;Y_{1})+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}),I(U,W;Y_{1},\hat{Y}_{2})\right\}$
$\qquad+I(V,W;Y_{2})+C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})-I(U;V|W)$ (27)
Remark 2.
It is clear that an alternative achievable rate region can be derived by exchanging the cooperative protocol at the receivers in the achievability scheme of Proposition 1, i.e., to apply quantize-bin-and-forward at the receiver $Y_{1}$ first and then a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{2}$.
As discussed before, the parameter $\alpha_{1}$ in the achievable rate region $\mathcal{R}_{i}^{(1)}$ reflects a balance between the rates associated with the decode-and-forward and the quantize-bin-and-forward schemes at the receiver that applies both of them (i.e., $Y_{1}$). It turns out that the optimal value for the parameter $\alpha_{1}$ can be determined. This results in a simplified expression for the achievable rate region as given in the following theorem.
Theorem 2.
For the two-user BC with conferencing decoders shown in Fig. 2, the rate region expressed in (21)-(27) over some joint PDF $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$ is achievable. This achievable rate region is obtained by setting the parameter $\alpha_{1}$ in $\mathcal{R}_{i}^{(1)}$ optimally as $\alpha_{1}^{*}$, given below:
$\alpha_{1}^{*}=\min\left\{\frac{I(\hat{Y}_{1};U,Y_{1}|W,Y_{2})}{C_{12}},1\right\}.$ (26)
Proof:
First of all, $\zeta_{1}$ and $\zeta_{2}$ in the characterization of $\mathcal{R}_{i}^{(1)}$ can be respectively replaced by $\tilde{\zeta}_{1}$ and $\tilde{\zeta}_{2}$ given in (19)-(20). Now, by setting $\alpha_{1}=\alpha_{1}^{*}$ in (26), one can easily derive the characterization (21)-(27). To see that this choice is in fact optimal, it is sufficient to note that
$\min\Big{\{}I(V,W;Y_{2})+\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})+\bar{\alpha}_{1}C_{12},$
$\qquad\quad\qquad I(V,W;\hat{Y}_{1},Y_{2})+\bar{\alpha}_{1}C_{12}\Big{\}}$
$\qquad\leq I(V,W;Y_{2})+C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})$ (27)
and
$\min$$\left\{\begin{array}[]{l}I(V;Y_{2}|W)+\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2}),\\ I(V;\hat{Y}_{1},Y_{2}|W)\end{array}\right\}$ (30)
$\leq$ $\min\left\{\begin{array}[]{l}{I(V;Y_{2}|W)+C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2}),}\\ {I(V;\hat{Y}_{1},Y_{2}|W)}\end{array}\right\}.$ (33)
Remark 3.
The achievable rate region $\mathcal{R}_{i}^{(1)}$ given in Theorem 2 clearly includes both of the regions previously presented in [20, Theorem 2] and [23, Appendix B] as subsets. Moreover, after some algebraic computation, one can show that $\mathcal{R}_{i}^{(1)}$, restricted over the distributions of the form $P(u,v,w,x)P(\hat{y}_{1}|w,y_{1})P(\hat{y}_{2}|y_{2})$, includes the achievable rate region of [21, Theorem 3] as a subset. Therefore, $\mathcal{R}_{i}^{(1)}$ encompasses these previously known results.
Let us now specialize the achievable rate region $\mathcal{R}_{i}^{(1)}$ for a primitive relay channel [30]. The primitive relay channel is a communication scenario wherein a transmitter sends a message to a receiver with the help of a relay node which has a digital relay link of given capacity to the receiver. This scenario is a special case of the two-user BC with conferencing decoders by setting $M_{0}=M_{1}=\varnothing$ and $C_{21}=0$ in Fig. 2. Now, by setting $U=\varnothing$ and $V=X$ in $\mathcal{R}_{i}^{(1)}$, we obtain the following achievable rate for the primitive relay channel:
$\max_{P(w,x)P(\hat{y}_{1}|w,y_{1})}\min\Big{\{}{I(X;Y_{2})+C_{12}-I(\hat{Y}_{1};Y_{1}|X,W,Y_{2}),}$
$\qquad{I(W;Y_{1})+I(X;\hat{Y}_{1},Y_{2}|W),}$
$\qquad{I(W;Y_{1})+I(X;Y_{2}|W)+C_{12}-I(\hat{Y}_{1};Y_{1}|X,W,Y_{2})}\Big{\}}.$ (34)
This is in fact the largest known achievable rate for the primitive relay channel as discussed in [31].
$R_{0}+R_{1}$ $\leq I(U,W;Y_{1})+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$ (39)
$R_{0}+R_{2}$ $\leq I(V,W;Y_{2})$ (40)
$R_{0}+R_{1}+R_{2}$ $\leq\min\left\{\begin{array}[]{l}{I(U;Y_{1}|W)+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}),}\\ {I(U;Y_{1},\hat{Y}_{2}|W)}\end{array}\right\}+I(V,W;Y_{2})-I(U;V|W)$ (43)
$R_{0}+R_{1}+R_{2}$ $\leq I(U,W;Y_{1})+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$
$\qquad+\min\left\{\begin{array}[]{l}{I(V;Y_{2}|W)+C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2}),}\\ {I(V;\hat{Y}_{1},Y_{2}|W)}\end{array}\right\}-I(U;V|W)$ (46)
$2R_{0}+R_{1}+R_{2}$ $\leq I(U,W;Y_{1})+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})+I(V,W;Y_{2})-I(U;V|W)$ (47)
Next, we present a second achievable rate region based on a combination of decode-and-forward and quantize-bin-and-forward at one receiver first and then quantize-bin-and-forward at the other receiver as cooperation strategy. This strategy may be advantageous when part of the message is already decodable at the first receiver before conferencing.
Proposition 2.
Consider the two-user BC with conferencing decoders shown in Fig. 2. Let $\mathcal{R}_{i}^{(2)}$ denote the set of all rate triples $(R_{0},R_{1},R_{2})\in\mathbb{R}_{+}^{3}$ such that
$R_{0}+R_{1}$ $\leq\min\Big{\{}I(U,W;Y_{1})+\eta_{2}+\bar{\alpha}_{2}C_{21},$
$\qquad I(U,W;Y_{1},\hat{Y}_{2})+\bar{\alpha}_{2}C_{21}\Big{\}}$ (35)
$R_{0}+R_{2}$ $\leq I(V,W;Y_{2})$ (36)
$R_{0}+R_{1}+R_{2}$ $\leq\min\Big{\{}I(U;Y_{1}|W)+\eta_{2},I(U;Y_{1},\hat{Y}_{2}|W)\Big{\}}$
$\ +I(V,W;Y_{2})-I(U;V|W)$ (37)
$R_{0}+R_{1}+R_{2}$ $\leq\min\Big{\{}I(U,W;Y_{1})+\eta_{2}+\bar{\alpha}_{2}C_{21},$
$\qquad\qquad I(U,W;Y_{1},\hat{Y}_{2})+\bar{\alpha}_{2}C_{21}\Big{\}}$
$\ +\min\left\{I(V;Y_{2}|W)+\eta_{1},I(V;\hat{Y}_{1},Y_{2}|W)\right\}$
$\ -I(U;V|W)$ (38)
$2R_{0}+R_{1}+R_{2}$ $\leq\min\Big{\{}I(U,W;Y_{1})+\eta_{2}+\bar{\alpha}_{2}C_{21},$
$\qquad\qquad I(U,W;Y_{1},\hat{Y}_{2})+\bar{\alpha}_{2}C_{21}\Big{\}}$
$\ +I(V,W;Y_{2})-I(U;V|W)$ (39)
where
$\eta_{1}$ =$=\{C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (40)
$\eta_{2}$ =$=\{\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (41)
for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|w,y_{2})$ and $\alpha_{2}\in[0,1]$. The convex closure of the set $\mathcal{R}_{i}^{(2)}$ is achievable.
Proof:
The cooperation strategy to achieve the rate region $\mathcal{R}_{i}^{(2)}$ is to apply Marton’s coding with a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{2}$ first and then quantize-bin-and-forward at the receiver $Y_{1}$. At the receiver $Y_{2}$, the capacity of the conferencing link $C_{21}$ is divided into two parts. A fraction $\bar{\alpha}_{2}C_{21}$ is devoted for decode-and-forward and the remaining $\alpha_{2}C_{21}$ is devoted for quantize-bin-and-forward scheme. A complete proof of Proposition 2 can be found in Appendix C. ∎
Remark 4.
In the characterization of the rate region (2)-(39), one can replace $\eta_{1}$ and $\eta_{2}$ with $\tilde{\eta}_{1}$ and $\tilde{\eta_{2}}$, respectively, which are given as follows:
$\tilde{\eta}_{1}$ $=C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})$ (42)
$\tilde{\eta_{2}}$ $=\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}).$ (43)
To see why this is the case, it is sufficient to note that in the case of $C_{12}<I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})$, we can set $\hat{Y}_{1}=\varnothing$ and obtain the same rate region $\mathcal{R}_{i}^{(2)}$. Similarly, in the case of $\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, we can set $\hat{Y}_{2}=\varnothing$ and obtain the same rate region.
Remark 5.
One can obtain an alternative achievable rate region by exchanging the cooperative protocol at the receivers in the acheivability scheme of Proposition 2, i.e., to apply a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{1}$ first and then quantize-bin-and-forward at the receiver $Y_{2}$.
Similar to the region $\mathcal{R}_{i}^{(1)}$, we can derive the optimal value of the parameter $\alpha_{2}$ for the region $\mathcal{R}_{i}^{(2)}$.
Theorem 3.
For the two-user BC with conferencing decoders shown in Fig. 2, the rate region expressed in (39)-(47) over some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|w,y_{2})$ is achievable. This achievable rate region is obtained by setting the parameter $\alpha_{2}$ in $\mathcal{R}_{i}^{(2)}$ optimally as $\alpha_{2}^{*}$, given below:
$\alpha_{2}^{*}=\min\left\{\frac{I(\hat{Y}_{2};Y_{2}|W,Y_{1})}{C_{21}},1\right\}.$ (49)
Proof:
First of all, $\eta_{1}$ and $\eta_{2}$ in the characterization (2)-(39) are respectively replaced by $\tilde{\eta}_{1}$ and $\tilde{\eta}_{2}$ given in (42). Now, by setting $\alpha_{2}=\alpha_{2}^{*}$ in (49), we easily obtain the characterization (39)-(47). To prove that this choice is in fact optimal, it is sufficient to consider the following inequalities:
$\min\Big{\{}I(U,W;Y_{1})+\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})+\bar{\alpha}_{2}C_{21},$
$\qquad\qquad I(U,W;Y_{1},\hat{Y}_{2})+\bar{\alpha}_{2}C_{21}\Big{\}}$
$\qquad\leq I(U,W;Y_{1})+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$ (50)
and
$\min\left\{\begin{array}[]{l}I(U;Y_{1}|W)+\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}),\\ I(U;Y_{1},\hat{Y}_{2}|W)\end{array}\right\}$ (53)
$\leq\min\left\{\begin{array}[]{l}{I(U;Y_{1}|W)+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}),}\\ {I(U;Y_{1},\hat{Y}_{2}|W)}\end{array}\right\}$ (56)
and by noting that $I(\hat{Y}_{2};U,Y_{2}|W,Y_{1})=I(\hat{Y}_{2};Y_{2}|W,Y_{1})$. ∎
In general, it is not easy to compare the two regions $\mathcal{R}_{i}^{(1)}$ and $\mathcal{R}_{i}^{(2)}$ numerically. An advantage of the region $\mathcal{R}_{i}^{(2)}$ over the region $\mathcal{R}_{i}^{(1)}$ is that its feasible set of distributions is a larger set. We further note that by allowing a fourth auxiliary random variable, it is possible to design an achievability scheme that includes both of $\mathcal{R}_{i}^{(1)}$ and $\mathcal{R}_{i}^{(2)}$ as special cases. However, evaluating the resulting rate region would be complex.
III-C Capacity Regions for Specific Channels
The achievability results in the previous section may be further improved, e.g., by applying multiple rounds of cooperation at the receivers. However, in what follows, we demonstrate that the rate region $\mathcal{R}_{i}^{(1)}$ is already sufficient to establish new capacity results for several particular two-user BCs with conferencing decoders (as well as approximate capacity results for the Gaussian channel as seen in the next section). The first such capacity result is for a class of semi-deterministic BCs with degraded message sets. Note that the determinism here is more general than that of semi-deterministic BC in [23].
Theorem 4.
Consider the two-user BC with degraded message sets, with a common message for both users and a private message for the first user, and bidirectional conferencing decoders. If the BC satisfies certain semi-deterministic property in the sense that $Y_{2}=f(X,Y_{1})$, then its capacity region is given by the closure of the convex hull of all $(R_{0},R_{1})$ satisfying
$R_{0}$ $\leq I(V;Y_{2})+C_{12}$ (57)
$R_{0}+R_{1}$ $\leq I(X;Y_{1})+C_{21}$ (58)
$R_{0}+R_{1}$ $\leq I(X;Y_{1}|V)+I(V;Y_{2})+C_{12}+C_{21}$ (59)
$R_{0}+R_{1}$ $\leq I(X;Y_{1},Y_{2}|V)+I(V;Y_{2})+C_{12}$ (60)
$R_{0}+R_{1}$ $\leq I(X;Y_{1},Y_{2})$ (61)
for some joint PDFs $P(v,x)$, where $V\rightarrow X\rightarrow Y_{1},Y_{2}$ forms a Markov chain and $|\mathcal{V}|\leq|\mathcal{X}|+2$.
Proof:
The achievability is derived by setting $U=X$, $W=V$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$. The converse proof is readily given by $\mathcal{R}_{o}$. ∎
Note that $(Y_{1},Y_{2})$ appears jointly in the first mutual information expression in (60). Thus, the capacity region of this BC depends on the joint distribution $p(y_{1},y_{2}|x)$ and not just on the marginals, in contrast to BC without conferencing decoders.
Remark 6.
By setting $R_{0}=0$ and $C_{12}=0$, the result of Theorem 4 is reduced to the capacity of deterministic primitive relay channel derived in [32]. In this case, the capacity is given by $\max_{P(x)}\min\{I(X;Y_{1})+C_{21},I(X;Y_{1},Y_{2})\}$. Note that for the case of $R_{0}=0$ and $C_{12}=0$, it is optimal to set $V=\varnothing$. In fact, the cut-set bound is achievable. This capacity is achieved by quantize-bin-and-forward as the relay strategy.
We present two interesting examples of this class of BC.
Example 1.
Consider the following binary channel,
$\begin{split}\begin{cases}Y_{1}=X\oplus Z\\ Y_{2}=Z\end{cases}\end{split}$ (62)
where $Z$ is a binary noise and $\oplus$ is the XOR operator. For this channel, we have $Y_{2}=X\oplus Y_{1}$, so the channel has the desired deterministic property. The capacity region of this BC with degraded message sets is given by all $(R_{0},R_{1})$ satisfying
$R_{0}$ $\leq C_{12}$ (63)
$R_{0}+R_{1}$ $\leq I(X;Y_{1})+C_{21}$ (64)
$R_{0}+R_{1}$ $\leq H(X)$ (65)
for some $P(x)$.
In this example, the receiver $Y_{2}$ does not receive information from the transmitter directly. Instead, it observes the noise $Z$ and relays the noise to the receiver $Y_{1}$ by sending a compressed version of it through the digital link $C_{21}$. The receiver $Y_{1}$ then decodes both messages and forwards the common message to the receiver $Y_{2}$ through the digital link $C_{12}$. Therefore, the receiver $Y_{2}$ can still receive information at a positive rate.
Example 2.
Consider the following binary channel,
$\begin{cases}Y_{1}=Z\\ Y_{2}=X\oplus Z\end{cases}$ (66)
where again $Z$ is a binary noise. The capacity region of this BC with degraded message sets is the closure of the convex hull of all $(R_{0},R_{1})$ satisfying
$R_{0}$ $\leq I(V;Y_{2})+C_{12}$ (67)
$R_{0}+R_{1}$ $\leq C_{21}$ (68)
$R_{0}+R_{1}$ $\leq H(X|V)+I(V;Y_{2})+C_{12}$ (69)
$R_{0}+R_{1}$ $\leq H(X)$ (70)
for some $P(v,x)$, where $V\rightarrow X\rightarrow Y_{1},Y_{2}$ forms a Markov chain and $|\mathcal{V}|\leq 3$.
Figure 3: Capacity region and cut-set outer bound for the BC in Example 2, where $Z$ is a binary noise with ${P(Z=0)=0.2}$.
Unlike the previous example, where the capacity region coincides with the cut-set bound, the capacity region for this BC is characterized using an auxiliary variable $V$. In Fig. 3, we plot the capacity region along with the cut-set outer bound for the channel in Example 2 with $P(Z=0)=0.2$ for two cases: $(C_{12},C_{21})=(0,0.9)$ and $(C_{12},C_{21})=(0.1,0.9)$, by searching over all possible $P(v,x)$. As can be seen, the capacity region lies strictly below the cut-set bound in this example. Moreover, $C_{12}$ strictly improves the capacity region, although $Y_{1}$ only observes the noise. In fact, the constraint (69) (which arises from (60)) plays a crucial role in characterizing the capacity region. Without (69), the remaining constraints simply reduce to the cut-set bound.
For this channel, the optimal coding strategy is as follows. Note that the receiver $Y_{1}$ (which is supposed to decode both common and private messages) does not receive any information from the transmitter directly. The optimal cooperation strategy is that the receiver $Y_{2}$ first sends a compressed version of its received signal to the receiver $Y_{1}$ through the digital link $C_{21}$. Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal which is in fact the channel noise) and then forwards part of the common message (using the variable $V$) to the receiver $Y_{2}$ through the link $C_{12}$. Lastly, the receiver $Y_{2}$ decodes the common message using its received signal.
It is worthwhile to make the following observation. For Example 2, as the receiver $Y_{1}$ observes the channel noise only, one might think that a cooperative scheme in which $Y_{1}$ applies quantize-bin-and-forward and $Y_{2}$ applies decode-and-forward is the right strategy. However, it turns out that such a scheme is not optimal when the message to $Y_{2}$ is a degraded version of the message to $Y_{1}$. This is in contrast to Example 1, in which the message set degrades in the opposite direction and performing quantize-bin-and-forward at the receiver that observes the noise only is the capacity-achieving strategy.
As far as we know, the above results are the first cases where a combination of quantize-bin-and-forward and decode-and-forward strategies yields an optimal bidirectional cooperation protocol. There is no previously known capacity result in the literature for channels with bidirectional cooperation between the receiver. Moreover, these results demonstrate that a one-round cooperation scheme is sufficient for achieving capacity for this class of channels.
It is also possible to provide a capacity result for a new class of more capable semi-deterministic BCs with both common and private messages and one-sided conferencing.
Theorem 5.
Consider the two-user BC with both common and private messages and with certain semi-deterministic property in the sense that $Y_{2}=f(X,Y_{1})$. Moreover, assume that the channel is more-capable, i.e., $I(X;Y_{2})\leq I(X;Y_{1})$ for every input distribution $P(x)$. For the channel with one-sided conferencing, i.e., with a digital relay link of finite capacity $C_{21}$ from the decoder at $Y_{2}$ to the decoder at $Y_{1}$, the capacity region is given by the closure of the convex hull of all $(R_{0},R_{1},R_{2})$ satisfying
$R_{0}+R_{2}$ $\leq I(V;Y_{2})$ (71)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{1})+C_{21}$ (72)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{1}|V)+I(V;Y_{2})+C_{21}$ (73)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{1},Y_{2}|V)+I(V;Y_{2})$ (74)
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{1},Y_{2})$ (75)
for some joint PDFs $P(v,x)$, where $V\rightarrow X\rightarrow Y_{1},Y_{2}$ forms a Markov chain and $|\mathcal{V}|\leq|\mathcal{X}|+2$.
Proof:
The achievability is derived by setting $U=X$, $W=V$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$.
For the converse, note that due to the Markov chain condition $U\rightarrow X\rightarrow Y_{1},Y_{2}$, the more-capable condition implies that
$I(X;Y_{2}|U)\leq I(X;Y_{1}|U)\quad\forall P(u,x).$ (76)
Therefore, the constraint (8) from the outer bound $\mathcal{R}_{o}$ with $C_{12}=0$ becomes
$R_{0}+R_{1}+R_{2}$ $\leq I(X;Y_{2}|U)+I(U;Y_{1})+C_{21}$ (77)
$\leq I(X;Y_{1}|U)+I(U;Y_{1})+C_{21}$ (78)
$=I(X;Y_{1})+C_{21},$ (79)
where the last step is again due to the Markov chain. The remaining constraints are directly derived from $\mathcal{R}_{o}$ by setting $C_{12}=0$. The resulting outer bound coincides with $\mathcal{R}_{i}^{(1)}$. ∎
Note that the structure of the novel outer bound $\mathcal{R}_{o}$ given in Theorem 1, in particular, with respect to the constraints (9) and (10), is crucial for deriving the capacity results in Theorems 4 and 5. In the next section, we show that the bounds can also be used to derive capacity results for certain classes of Gaussian BCs.
IV Gaussian BC with Conferencing Decoders
We now study the Gaussian BC with conferencing decoders and with correlated noises as shown in Fig. 4. The channel model is defined as follows:
$\begin{cases}Y_{1}=aX+Z_{1}\\ Y_{2}=bX+Z_{2}\end{cases}$ (80)
where $Z_{1}$ and $Z_{2}$ are correlated Gaussian noises with zero mean and unit variance and correlation coefficient $\lambda$, i.e., $\mathbb{E}[Z_{1}Z_{2}]=\lambda$, $X$ is the input signal with $\mathbb{E}[X^{2}]\leq P$, and $a$, and $b$ are the real-valued channel gains. Throughout this section, the transmit power constraint is assumed to be strictly positive, i.e., $P>0$. The Gaussian BC is equipped with a conferencing link of capacity $C_{12}$ from the decoder at $Y_{1}$ to the decoder at $Y_{2}$, and a conference link of capacity $C_{21}$ from the decoder at $Y_{2}$ to the decoder at $Y_{1}$.
We show that the bounds derived in earlier part of the paper are useful for deriving capacity and approximate capacity results for various types of Gaussian BC with conferencing decoders.
Figure 4: Two-user Gaussian BC with power constraint $P>0$ and noise correlation $\lambda$, and with conferencing links of capacities $C_{12}$ and $C_{21}$.
IV-A Converse
We begin with the outer bound. By taking into consideration the input power constraint $\mathbb{E}[X^{2}]\leq P$, the outer bound in Theorem 1 can be optimized over its auxiliary variables $U$ and $V$ for the Gaussian channel. This gives an explicit characterization of the mutual information terms in the outer bound as derived below.
Theorem 6.
Consider the Gaussian BC (80) with conferencing decoders. Assume that $|a|\geq|b|$. Let $\mathcal{R}_{o}^{G}$ denote the set of all rate triples $(R_{0},R_{1},R_{2})$ such that for some $\alpha,\beta\in[0,1]$
$R_{0}+R_{1}$ $\leq\psi\left(\frac{(1-\alpha)a^{2}P}{\alpha a^{2}P+1}\right)+C_{21}$ (81)
$R_{1}$ $\leq\Psi_{2}+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$ (82)
$R_{0}+R_{2}$ $\leq\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{12}$ (83)
$R_{2}$ $\leq\Psi_{1}+\psi\left(\frac{(1-\alpha)b^{2}P}{\alpha b^{2}P+1}\right)$ (84)
$R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\beta a^{2}P\right)+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$
$\qquad\qquad\qquad\qquad+C_{12}+C_{21}$ (85)
$R_{0}+R_{1}+R_{2}$ $\leq\Psi_{2}+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{12}$ (86)
$R_{0}+R_{1}+R_{2}$ $\leq\Psi_{1}+\psi\left(\frac{(1-\alpha)a^{2}P}{\alpha a^{2}P+1}\right)+C_{21}$ (87)
$R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab}{1-\lambda^{2}}\right)P\right),$ (88)
where
$\Psi_{1}$ $=\psi\left(\alpha\left(\frac{a^{2}+b^{2}-2\lambda ab}{1-\lambda^{2}}\right)P\right)$ (89)
$\Psi_{2}$ $=\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab}{1-\lambda^{2}}\right)P\right).$ (90)
The set $\mathcal{R}_{o}^{G}$ constitutes an outer bound on the capacity region of the BC.
Proof:
The outer bound is based on applying the entropy power inequality to the outer bound derived in Theorem 1. The details are given in Appendix D. ∎
Remark 7.
For the Gaussian BC (80) with $|a|\geq|b|$, when $\lambda=\frac{b}{a}$, we have
$\frac{a^{2}+b^{2}-2\lambda ab}{1-\lambda^{2}}=a^{2}.$ (91)
Note that in the case of $\lambda=1$ (with $a=b$), or $\lambda=-1$ (with $a=-b$), the above equality is to be interpreted in the limiting sense. Further, when $\lambda=\frac{b}{a}$, the BC is physically degraded, for which the outer bound $\mathcal{R}_{o}^{G}$ is achievable. (See Remark 11.)
We now discuss the BC with perfectly correlated noises, i.e., $\lambda=1$ or $-1$, but with the assumption that $\lambda\neq\frac{b}{a}$.
IV-B Capacity Regions for Specific Channels
Theorem 7.
Consider the two-user Gaussian BC (80) with degraded message sets (i.e., a common message for both receivers at rate $R_{0}$ and a private message for the first receiver at rate $R_{1}$) and bidirectional conferencing decoders. Assume $P>0$. When the noises are perfectly correlated, i.e., $|\lambda|=1$, but $\lambda\neq\frac{b}{a}$, the capacity region of the channel is given by the closure of the convex hull of all $(R_{0},R_{1})$ satisfying
$R_{0}$ $\leq\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{12}$ (92)
$R_{0}+R_{1}$ $\leq\psi\left(a^{2}P\right)+C_{21}$ (93)
$R_{0}+R_{1}$ $\leq\psi\left(\beta a^{2}P\right)+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{12}+C_{21}$ (94)
for some $\beta\in[0,1]$.
Proof:
The achievability is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}$. Note that by this choice of random variables, when $|\lambda|=1$, we have
$I(\hat{Y}_{2};Y_{2}|W,U,Y_{1})=\frac{1}{2}\psi\left(\frac{1-\lambda^{2}}{\hat{\sigma}^{2}}\right)=0$ (95)
Moreover,
$I(X;Y_{1},\hat{Y}_{2})$ $=\frac{1}{2}\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
$=\frac{1}{2}\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right)$ (96)
and
$I(X;Y_{1},\hat{Y_{2}}|W)$ $=\frac{1}{2}\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
$=\frac{1}{2}\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right).$ (97)
Note that if $|\lambda|=1$ but $\lambda\neq\frac{b}{a}$, then $a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}\neq 0$. Therefore, by letting $\hat{\sigma}^{2}\rightarrow 0$, one can make the above two mutual information terms arbitrarily large, so they are would not be in effect in the characterization of the achievable rate region. For the case of $|a|\geq|b|$, the converse proof is readily given by $\mathcal{R}_{o}^{G}$ in Theorem 6. For the case of $|a|<|b|$, the achievable rate region (92)-(94) is maximized for $\beta=0$ and it coincides with the cut-set bound. ∎
The capacity result in Theorem 7 is among the rare cases in network information theory for which the quantize-bin-and-forward strategy contributes to achieving capacity.
Example 3.
A special case of the Gaussian BC with perfectly correlated noises is the following channel:
$\begin{cases}Y_{1}=X+Z\\ Y_{2}=X-Z\end{cases}$ (98)
where $Z$ is a zero mean unit variance Gaussian noise. In this scenario, the two receivers see exactly the same noise but with a different sign. For this channel, the capacity result of Theorem 7 reduces to the following:
$R_{0}$ $\leq\psi(P)+C_{12}$ (99)
$R_{0}+R_{1}$ $\leq\psi(P)+C_{21}.$ (100)
In fact for this channel, the cut-set bound is achievable. To achieve this capacity, the second receiver applies quantize-bin-and-forward and the first receiver applies decode-and-forward as the cooperation protocol.
Observe that in the case of common message only, even as $P\rightarrow 0$, it is possible to achieve a strictly positive rate of $R_{0}=\min\{C_{12},C_{21}\}$. This makes concrete the statement made in the introduction, namely, that it is possible to use an infinitesimally small amount of power to transmit information at a strictly positive rate. Note that the argument presented in the proof of Theorem 7 is only valid for strictly positive values of input power $P$. We summarize this observation in the following remark.
Remark 8.
The capacity characterization given in Theorem 7 gives rise to the following interesting observation. Even with a very small (yet positive) amount of input power $P$, one can transmit information over the channel at a rate as high as the capacities of the conferencing links, if the noises are perfectly correlated, i.e., $|\lambda|=1$, (assuming $\lambda\neq\frac{b}{a}$). In fact, as $P\rightarrow 0$, the capacity region of the Gaussian BC with degraded message sets is as follows:
$R_{0}$ $\leq\epsilon_{1}(P)+C_{12}$ (101)
$R_{0}+R_{1}$ $\leq\epsilon_{2}(P)+C_{21}$ (102)
for some $\epsilon_{1}(P)$ and $\epsilon_{2}(P)$, which go to zero as $P\rightarrow 0$.
Next, we draw a connection between the BC with conferencing decoders and the relay channel.
Remark 9.
For the special case of primitive relay channel with perfectly correlated noise, where $Y_{2}$ acts as a relay for $Y_{1}$, $C_{12}=0$ and there is only a private message for the first receiver, the capacity result given in Theorem 7 is reduced to $R_{1}\leq\psi\left(a^{2}P\right)+C_{21}$. In this case, using the quantize-bin-and-forward strategy, one can achieve the cut-set bound. Interestingly, the capacity does not depend on $b$ at all.
This result is an example of a semi-deterministic primitive relay channel, because the relay observation $Y_{2}=bX+Z_{2}$ is a deterministic function of the input $X$ and the receiver observation $Y_{1}=aX+Z_{1}$, when $Z_{1}$ and $Z_{2}$ are perfectly correlated and $\lambda\neq\frac{b}{a}$. In this case, the cut-set bound is achievable [32].
The following two examples further illustrate Theorem 7.
Example 4.
Consider the following Gaussian BC:
$\begin{cases}Y_{1}=X+Z\\ Y_{2}=Z\end{cases}$ (103)
where $Z$ is a zero mean unit variance Gaussian noise. In this case, the receiver $Y_{2}$ does not receive any information from the transmitter and only observes the additive noise of the receiver $Y_{1}$. For this channel, the capacity region can be derived by setting $a=1$ and $b=0$ in (92)-(94) and is given by
$R_{0}$ $\leq C_{12}$ (104)
$R_{0}+R_{1}$ $\leq\psi(P)+C_{21}.$ (105)
This is the Gaussian counterpart of Example 1. To achieve this capacity, first the receiver $Y_{2}$ sends a compressed version of the observed noise $Z$ to the receiver $Y_{1}$ through the digital link $C_{21}$. Next, the receiver $Y_{1}$ decodes both common and private messages based on the message from $Y_{2}$, then forwards the common message to $Y_{2}$ through the digital link $C_{12}$.
Example 5.
Consider the following Gaussian BC:
$\begin{cases}Y_{1}=Z\\ Y_{2}=X+Z.\end{cases}$ (106)
In this case, the capacity region is given by
$R_{0}$ $\leq\psi(P)+C_{12}$ (107)
$R_{0}+R_{1}$ $\leq C_{21}.$ (108)
This is the Gaussian counterpart of Example 2. Similar to Example 2, the receiver $Y_{1}$ (which needs to detect both common and private messages) does not receive any information from the transmitter directly. The capacity achieving cooperation protocol is that the receiver $Y_{2}$ first sends a compressed version of its received signal to the receiver $Y_{1}$ through the digital link $C_{21}$. Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal, which is in fact just the channel noise) then forwards the bin index of the decoded common message to the receiver $Y_{2}$ through the link $C_{12}$. Lastly, the receiver $Y_{2}$ decodes the common message using the bin index received and its received signal.
Note that in Example 5, as the receiver $Y_{1}$ observes the channel noise only, one might think that a cooperative scheme in which $Y_{1}$ applies quantize-bin-and-forward first followed by $Y_{2}$ applying decode-and-forward should be used. However, such a scheme is not optimal in this particular BC where the message for $Y_{2}$ is a degraded version of the message for $Y_{1}$.
Theorem 7 is a capacity result for the Gaussian BC with degraded message sets, and with bidirectional conferencing decoders and perfectly correlated noises. Next, we establish the capacity region of a Gaussian BC with both common and private messages, and with one-sided cooperation.
Theorem 8.
Consider the two-user Gaussian BC (80) with both common and private messages where $|a|\geq|b|$ and only the weaker receiver is connected to the stronger one by a conferencing link, i.e., $C_{12}=0$. Assume $P>0$. For the channel with perfectly correlated noises with $|\lambda|=1$, but $\lambda\neq\frac{b}{a}$, the capacity region is given by the closure of the convex hull of all $(R_{0},R_{1},R_{2})$ satisfying
$R_{0}+R_{2}$ $\leq\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$ (109)
$R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\beta a^{2}P\right)+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{21}$ (110)
for some $\beta\in[0,1]$.
The proof of Theorem 8 is similar to that of Theorem 7 and therefore omitted.
For the Gaussian BC in which the noises are not perfectly correlated, the inner and outer bounds of this paper can provide approximate capacity results. The first such result is for the Gaussian BC with degraded message sets.
Theorem 9.
Consider the two-user Gaussian BC (80) with degraded message sets (i.e., a common message for both receivers and a private message for the first receiver) and bidirectional conferencing decoders. Assume that $|a|\geq|b|$. For all channel parameters $a$, $b$, $C_{12}$, $C_{21}$, and $\lambda$ with $|\lambda|<1$, the following achievable rate region is within half a bit from the capacity region:
$R_{0}$ $\leq\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{12}$ (111)
$R_{0}+R_{1}$ $\leq\psi\left(a^{2}P\right)+\{C_{21}-1/2\}^{+}$ (112)
$R_{0}+R_{1}$ $\leq\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+(1-\lambda^{2})a^{2}}{2(1-\lambda^{2})}\right)P\right)$ (113)
$R_{0}+R_{1}$ $\leq\psi\left(\beta a^{2}P\right)+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$
+$\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}+C_{12}$ (114)
$R_{0}+R_{1}$ $\leq\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+(1-\lambda^{2})a^{2}}{2(1-\lambda^{2})}\right)P\right)$
$\qquad\qquad\qquad+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{12}$ (115)
for some $\beta\in[0,1]$.
Proof:
The above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
By a comparison, one can see that the right-hand sides of the constraints (111), (112), (113), (114) and (115) are within half a bit of (83), (81), (88), (85), and (86), respectively. ∎
Next, we present an approximate capacity result for the Gaussian BC with one-way conferencing.
Theorem 10.
Consider the two-user Gaussian BC (80) with both common and private messages, where $|a|\geq|b|$ and only the weaker receiver has a conferencing link to the stronger receiver, i.e., $C_{12}=0$. For all channel parameters $a$, $b$, $C_{21}$, and $\lambda$ with $|\lambda|<1$, the following achievable rate region is within half a bit from the capacity region:
$R_{0}+R_{2}$ $\leq\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$ (116)
$R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+(1-\lambda^{2})a^{2}}{2(1-\lambda^{2})}\right)P\right)$ (117)
$R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\beta a^{2}P\right)+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$
+$\qquad\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}$ (118)
$R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+(1-\lambda^{2})a^{2}}{2(1-\lambda^{2})}\right)P\right)$
$\qquad\qquad\qquad+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$ (119)
for some $\beta\in[0,1]$.
Proof:
Similar to Theorem 9, the above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
By a simple comparison, one can see that the right-hand sides of the constraints (116), (117), (118) and (119) are within half a bit of (83), (88), (85), and (86), respectively. ∎
Finally, we derive an approximate capacity result for the two-user Gaussian BC (80) with both common and private messages and with bidirectional cooperative receivers. First, we present an achievable rate region for the channel using only one-way conferencing with decode-and-forward. The other conferencing link is not used, so the resulting achievable rate region is a sub-region of $\mathcal{R}_{i}^{(1)}$ (21)-(27). It turns out that this region is already approximately optimal when the noise correlation is small.
Proposition 3.
Let $\mathcal{R}_{i}^{DF-G}$ denote the set of all rate triples $(R_{0},R_{1},R_{2})$ such that
$R_{0}+R_{2}$ $\leq\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{12}$ (120)
$R_{0}+R_{1}+R_{2}$ $\leq\psi\left(a^{2}P\right)$ (121)
$R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\beta a^{2}P\right)+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)+C_{12}$ (122)
for some $\beta\in[0,1]$. The set $\mathcal{R}_{i}^{DF-G}$ constitutes an inner bound on the capacity region of the Gaussian BC (80) with conferencing decoders.
Proof:
The bound $\mathcal{R}_{i}^{DF-G}$ is derived from $\mathcal{R}_{i}^{(1)}$ given in Theorem 2 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{2}=\varnothing$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $\beta P$ and $(1-\beta)P$, respectively. Note that this inner bound is in fact derived for the channel with one-sided cooperation (i.e., it does not make use of the conferencing link $C_{21}$) using the decode-and-forward strategy alone. Nevertheless, it is a valid inner bound for the channel with bidirectional cooperation. ∎
Theorem 11.
Consider the two-user Gaussian BC (80) with both common and private messages and with bidirectional conferencing decoders. Assume that $|a|\geq|b|$. For all channel parameters $a$, $b$, $C_{12}$, $C_{21}$, and $\lambda$ with $|\lambda|<1$, the inner bound $\mathcal{R}_{i}^{DF-G}$ is within $\frac{1}{2}\log(\frac{2}{1-\left|\lambda\right|})$ bits of the capacity region.
Proof:
The constraint (120) is identical to (83). Moreover, by simple algebraic computation, one can show that the right-hand sides of the constraints (121) and (122) are within $\frac{1}{2}\log(\frac{2}{1-\left|\lambda\right|})$ bits of (88) and (86), respectively. ∎
Remark 10.
For the Gaussian BC with $\lambda ab\geq 0$, a better approximate capacity gap bound of $\frac{1}{2}\log(\frac{2}{1-\lambda^{2}})$ bits is possible for the region $\mathcal{R}_{i}^{DF-G}$ given in Proposition 3.
Theorem 11 states that if we do not use the quantize-bin-and-forward part of the conferencing protocol and rely solely on decode-and-forward, the gap to capacity would depend on the noise correlation. This is because decode-and-forward cannot exploit the noise correlation. Nevertheless, there is one special case for which decode-and-forward is optimal.
Remark 11.
For the Gaussian BC (80) with $|a|\geq|b|$, when $\lambda=\frac{b}{a}$, one can verify that the BC is physically degraded (see Appendix D) and the decode-and-forward achievable region $\mathcal{R}_{i}^{DF-G}$ of Proposition 3 coincides with the outer bound $\mathcal{R}_{o}^{G}$ in Theorem 6; (see (91)). Thus, $\mathcal{R}_{i}^{DF-G}$ is the capacity region.
As a concluding remark for this section, the novel structure of the outer bound given in Theorem 6, in particular, the constraint (86), is crucial for deriving the exact capacity results in Theorems 7 and 8, and also the approximate capacity results in Theorems 9, 10, and 11.
V Conclusions
Receiver cooperation can significantly improve the capacity of a broadcast channel. In this paper, we establish novel outer bounds, which are tighter than the cut-set bound, for the BC with bidirectional conferencing decoders using the Csiszár-Körner identity. Together with the achievability results based on Marton’s coding and judicious combination of quantize-bin-and-forward and decode-and-forward, we derive capacity results for specific classes of semi-deterministic BCs with degraded message sets and more capable semi-deterministic BCs with one-sided conferencing. Furthermore, for the Gaussian BC, we illustrate the importance of noise correlation when the receivers can cooperate using the conferencing links. For a Gaussian BC with perfectly negatively correlated noises, an infinitesimal amount of power is already sufficient for achieving a rate equal to the capacity of the conferencing links.
Appendix A Proof of Theorem 1
We only derive the constraints involving the auxiliary variable $U$, i.e., (1), (5), (6), (8), and (10). The constraints involving $V$ can be derived symmetrically. The last constraint is due to the cut-set bound.
Consider a length-$n$ code with common and private message rates $(R_{0},R_{1},R_{0})$ and with vanishing average probability of decoding error. Define the auxiliary variables as follows:
$U_{t}=(M_{0},M_{1},Y_{2}^{t-1},Y_{1,t+1}^{n}),\qquad t=1,...,n.$ (123)
By Fano’s inequality, we have
$n$ $(R_{0}+R_{1})$
$\leq I(M_{0},M_{1};Y_{1}^{n},J_{21}^{\Gamma})+n\epsilon_{n}^{1}$ (124)
$=I(M_{0},M_{1};Y_{1}^{n})+I(M_{0},M_{1};J_{21}^{\Gamma}|Y_{1}^{n})+n\epsilon_{n}^{1}$ (125)
$\leq\sum_{t=1}^{n}I(M_{0},M_{1};Y_{1,t}|Y_{1,t+1}^{n})+H(J_{21}^{\Gamma})+n\epsilon_{n}^{1}$ (126)
$\leq\sum_{t=1}^{n}I(M_{0},M_{1},Y_{1,t+1}^{n},Y_{2}^{t-1};Y_{1,t})+H(J_{21}^{\Gamma})+n\epsilon_{n}^{1}$ (127)
$\leq\sum_{t=1}^{n}I(U_{t};Y_{1,t})+nC_{21}+n\epsilon_{n}^{1}$ (128)
[…]
Next, we derive the constraints on $R_{2}$. By Fano’s inequality,
$nR_{2}$ $\leq I(M_{2};Y_{2}^{n},J_{12}^{\Gamma})+n\epsilon_{n}^{2}$ (129)
$\leq I(M_{2};Y_{1}^{n},Y_{2}^{n},J_{12}^{\Gamma})+n\epsilon_{n}^{2}$ (130)
$=I(M_{2};Y_{1}^{n},Y_{2}^{n})+n\epsilon_{n}^{2}$ (131)
$\leq I(M_{2};Y_{1}^{n},Y_{2}^{n}|M_{0},M_{1})+n\epsilon_{n}^{2}$ (132)
[…]
Now consider the following:
$I$ $(M_{2};Y_{1}^{n},Y_{2}^{n}|M_{0},M_{1})$
$=I(M_{2};Y_{2}^{n}|M_{0},M_{1})+I(M_{2};Y_{1}^{n}|Y_{2}^{n},M_{0},M_{1})$ (133)
$\leq I(M_{0},M_{1},M_{2};Y_{2}^{n})+I(M_{2};Y_{1}^{n}|Y_{2}^{n},M_{0},M_{1})$ (134)
$\leq\sum_{t=1}^{n}I(X_{t};Y_{2,t})$
$\qquad+\sum_{t=1}^{n}I(X_{t};Y_{1,t}|Y_{2,t},M_{0},M_{1},Y_{2}^{t-1},Y_{2,t+1}^{n},Y_{1,t+1}^{n})$ (135)
$=\sum_{t=1}^{n}I(X_{t};Y_{2,t})+\sum_{t=1}^{n}I(X_{t};Y_{1,t}|Y_{2,t},U_{t},Y_{2,t+1}^{n})$ (136)
$\leq\sum_{t=1}^{n}I(X_{t};Y_{2,t})+\sum_{t=1}^{n}I(X_{t};Y_{1,t}|Y_{2,t},U_{t}),$ (137)
[…]
In the same way, one can also derive
$I$ $(M_{2};Y_{1}^{n},Y_{2}^{n}|M_{0},M_{1})$
$=I(M_{2};Y_{1}^{n}|M_{0},M_{1})+I(M_{2};Y_{2}^{n}|Y_{1}^{n},M_{0},M_{1})$ (138)
$\leq\sum_{t=1}^{n}I(X_{t};Y_{1,t})+\sum_{t=1}^{n}I(X_{t};Y_{2,t}|Y_{1,t},U_{t}).$ (139)
By substituting (137) and (139) into (132), we obtain the desired bounds on $R_{2}$ (5) and (6).
Next, we derive the constraints on the sum rate:
$n$ $(R_{0}+R_{1}+R_{2})$
$\leq I(M_{0},M_{1};Y_{1}^{n},J_{21}^{\Gamma})+I(M_{2};Y_{2}^{n},J_{12}^{\Gamma})+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (140)
$\leq I(M_{0},M_{1};Y_{1}^{n})+I(M_{2};Y_{2}^{n})+I(M_{0},M_{1};J_{21}^{\Gamma}|Y_{1}^{n})$
$\qquad+I(M_{2};J_{12}^{\Gamma}|Y_{2}^{n})+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (141)
$\leq I(M_{0},M_{1};Y_{1}^{n})+I(M_{2};Y_{2}^{n}|M_{0},M_{1})$
$\qquad+H(J_{21}^{\Gamma})+H(J_{12}^{\Gamma})+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (142)
$\leq I(M_{0},M_{1};Y_{1}^{n})+I(M_{2};Y_{2}^{n}|M_{0},M_{1})$
$\qquad+n(C_{12}+C_{21})+n(\epsilon_{n}^{1}+\epsilon_{n}^{2}),$ (143)
where in (142) we use the fact that $M_{0}$, $M_{1}$ and $M_{2}$ are independent to conclude that $I(M_{2};Y_{2}^{n})\leq I(M_{2};Y_{2}^{n}|M_{0},M_{1})$.
Now consider the following:
$I$ $(M_{0},M_{1};Y_{1}^{n})+I(M_{2};Y_{2}^{n}|M_{0},M_{1})$
$=\sum_{t=1}^{n}I(M_{0},M_{1};Y_{1,t}|Y_{1,t+1}^{n})$
$\qquad+\sum_{t=1}^{n}I(M_{2};Y_{2,t}|M_{0},M_{1},Y_{2}^{t-1})$ (144)
$\leq\sum_{t=1}^{n}I(Y_{2}^{t-1},M_{0},M_{1};Y_{1,t}|Y_{1,t+1}^{n})$
$\qquad-\sum_{t=1}^{n}I(Y_{2}^{t-1};Y_{1,t}|M_{0},M_{1},Y_{1,t+1}^{n})$
$\qquad+\sum_{t=1}^{n}I(Y_{1,t+1}^{n},M_{2};Y_{2,t}|M_{0},M_{1},Y_{2}^{t-1})$ (145)
$=\sum_{t=1}^{n}I(Y_{2}^{t-1},M_{0},M_{1};Y_{1,t}|Y_{1,t+1}^{n})$
$\qquad-\sum_{t=1}^{n}I(Y_{2}^{t-1};Y_{1,t}|M_{0},M_{1},Y_{1,t+1}^{n})$
$\qquad+\sum_{t=1}^{n}I(Y_{1,t+1}^{n};Y_{2,t}|M_{0},M_{1},Y_{2}^{t-1})$
$\qquad+\sum_{t=1}^{n}I(M_{2};Y_{2,t}|M_{0},M_{1},Y_{2}^{t-1},Y_{1,t+1}^{n})$ (146)
$\leq\sum_{t=1}^{n}I(Y_{2}^{t-1},M_{0},M_{1};Y_{1,t}|Y_{1,t+1}^{n})$
$\qquad+\sum_{t=1}^{n}I(M_{2};Y_{2,t}|M_{0},M_{1},Y_{2}^{t-1},Y_{1,t+1}^{n})$ (147)
$\leq\sum_{t=1}^{n}I(Y_{2}^{t-1},Y_{1,t+1}^{n},M_{0},M_{1};Y_{1,t})$
$\qquad+\sum_{t=1}^{n}I(M_{2};Y_{2,t}|M_{0},M_{1},Y_{2}^{t-1},Y_{1,t+1}^{n})$ (148)
$\leq\sum_{t=1}^{n}I(U_{t};Y_{1,t})+\sum_{t=1}^{n}H(Y_{2,t}|U_{t})$
$\qquad-\sum_{t=1}^{n}H(Y_{2,t}|M_{0},M_{1},M_{2},Y_{2}^{t-1},Y_{1,t+1}^{n})$ (149)
$\leq\sum_{t=1}^{n}I(U_{t};Y_{1,t})+\sum_{t=1}^{n}H(Y_{2,t}|U_{t})$
$\qquad-\sum_{t=1}^{n}H(Y_{2,t}|X_{t},M_{0},M_{1},M_{2},Y_{2}^{t-1},Y_{1,t+1}^{n})$ (150)
$\leq\sum_{t=1}^{n}I(U_{t};Y_{1,t})+\sum_{t=1}^{n}H(Y_{2,t}|U_{t})$
$\qquad-\sum_{t=1}^{n}H(Y_{2,t}|X_{t},M_{0},M_{1},Y_{2}^{t-1},Y_{1,t+1}^{n})$ (151)
$\leq\sum_{t=1}^{n}I(U_{t};Y_{1,t})+\sum_{t=1}^{n}I(X_{t};Y_{2,t}|U_{t})$ (152)
where (147) is due to the Csiszár-Körner identity [27] [28, Lemma 7], which states that given two arbitrary random vectors $A_{1}^{n}$ and $B_{1}^{n}$ and an arbitrary random variable $C$, the following is true:
$\centering\sum_{t=1}^{n}I(A_{t+1}^{n};B_{t}|B_{1}^{t-1},C)=\sum_{t=1}^{n}I(B_{1}^{t-1};A_{t}|A_{t+1}^{n},C),\@add@centering$ (153)
and (149) is due to the definition of $U_{t}$, (150) is due to that conditioning does not increase entropy, (151) is due to the Markov chain $(M_{2},U_{t})\rightarrow X_{t}\rightarrow Y_{2,t}$. By substituting (152) in (143), we obtain the bound (8) on the sum rate.
Finally, we can write
$n$ $(R_{0}+R_{1}+R_{2})$
$\leq I(M_{0},M_{1};Y_{1}^{n},J_{21}^{\Gamma})+I(M_{2};Y_{2}^{n},J_{12}^{\Gamma})+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (154)
$\leq I(M_{0},M_{1};Y_{1}^{n})+I(M_{0},M_{1};J_{21}^{\Gamma}|Y_{1}^{n})$
$\qquad+I(M_{2};Y_{1}^{n},Y_{2}^{n},J_{12}^{\Gamma})+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (155)
$\leq I(M_{0},M_{1};Y_{1}^{n})+H(J_{21}^{\Gamma})+I(M_{2};Y_{1}^{n},Y_{2}^{n})$
$\qquad+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (156)
$\leq I(M_{0},M_{1};Y_{1}^{n})+nC_{21}+I(M_{2};Y_{1}^{n},Y_{2}^{n}|M_{0},M_{1})$
$\qquad+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (157)
$\leq I(M_{0},M_{1};Y_{1}^{n})+nC_{21}+I(M_{2};Y_{2}^{n}|M_{0},M_{1})$
$\qquad+I(M_{2};Y_{1}^{n}|Y_{2}^{n},M_{0},M_{1})+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (158)
$\leq\sum_{t=1}^{n}I(U_{t};Y_{1,t})+\sum_{t=1}^{n}I(X_{t};Y_{2,t}|U_{t})$
$\qquad+\sum_{t=1}^{n}I(X_{t};Y_{1,t}|Y_{2,t},U_{t})+nC_{21}+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (159)
$=\sum_{t=1}^{n}I(U_{t};Y_{1,t})+\sum_{t=1}^{n}I(X_{t};Y_{1,t},Y_{2,t}|U_{t})$
$\qquad+nC_{21}+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (160)
$=\sum_{t=1}^{n}I(U_{t};Y_{1,t})+\sum_{t=1}^{n}I(X_{t};Y_{1,t}|U_{t})$
$\qquad+\sum_{t=1}^{n}I(X_{t};Y_{2,t}|Y_{1,t},U_{t})+nC_{21}+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (161)
$=\sum_{t=1}^{n}I(U_{t},X_{t};Y_{1,t})+\sum_{t=1}^{n}I(X_{t};Y_{2,t}|Y_{1,t},U_{t})$
$\qquad+nC_{21}+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (162)
$=\sum_{t=1}^{n}I(X_{t};Y_{1,t})+\sum_{t=1}^{n}I(X_{t};Y_{2,t}|Y_{1,t},U_{t})$
$\qquad+nC_{21}+n(\epsilon_{n}^{1}+\epsilon_{n}^{2})$ (163)
[…]
Lastly, (11) is just the cut-set bound. By applying a standard time-sharing argument, the proof is thus complete. The cardinality bounds on $U$ and $V$ are due to Carathéodory theorem.
Appendix B Proof of Proposition 1
We prove the achievability of $\mathcal{R}_{i}^{(1)}$ using Marton’s coding for the BC together with quantize-bin-and-forward at $Y_{2}$, then a combination of decode-and-forward and quantize-bin-and-forward at $Y_{1}$. Fix a joint distribution
$P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$
and let $\alpha_{1}\in[0,1]$. We construct random codebooks consisting of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively.
First of all, each of the private messages $M_{1}$ and $M_{2}$ is split into two parts as follows:
$M_{1}$ $\longmapsto(M_{10},M_{11}),$ (164)
where $M_{10}\in[1:2^{nR_{10}}],M_{11}\in[1:2^{nR_{11}}]$, and
$M_{2}$ $\longmapsto(M_{20},M_{22}),$ (165)
where $M_{20}\in[1:2^{nR_{20}}],M_{22}\in[1:2^{nR_{22}}]$. Therefore,
$R_{1}=R_{10}+R_{11},$ (166)
$R_{2}=R_{20}+R_{22}.$ (167)
The sub-messages $M_{10}$ and $M_{20}$ are decoded at both receivers, i.e., the triple $(M_{0},M_{10},M_{20})$ is considered as the common message with rate
$R_{00}=R_{0}+R_{10}+R_{20}.$ (168)
Moreover, let $B_{1},B_{2},\hat{R}_{1},\hat{R}_{2}$ be non-negative real numbers denoting the rates of the binning steps and the compression steps in the proof. Also, define
$\hat{C}_{12}^{d}$ $=\min\{\bar{\alpha}_{1}C_{12},R_{00}\}$ (169)
$\hat{C}_{12}^{q}$ $=\min\{{\alpha}_{1}C_{12}\hat{R}_{1}\}$ (170)
$\hat{C}_{21}$ $=\min\{C_{21},\hat{R}_{2}\}$ (171)
as the portion of the conferencing link rates used in the two directions and for decode-and-forward and quantize-bin-and-forward, respectively.
B-A Encoding at the Transmitter
  1. 1.
    Generate at random $2^{nR_{00}}$ independent codewords $W^{n}$ according to $P(w^{n})=\prod_{t=1}^{n}P(w_{t})$. Label these codewords $W^{n}(m_{00})$ where $m_{00}\in[1:2^{nR_{00}}]$.
  2. 2.
    For each codeword $W^{n}(m_{00})$, generate at random $2^{n(R_{11}+B_{1})}$ independent codewords $U^{n}$ according to $\prod_{t=1}^{n}P(u_{t}|w_{t}(m_{00}))$. Label these codewords $U^{n}(m_{00},m_{11},b_{1})$ where $m_{11}\in[1:2^{nR_{11}}]$ and $b_{1}\in[1:2^{nB_{1}}]$.
  3. 3.
    For each codeword $W^{n}(m_{00})$, generate at random $2^{n(R_{22}+B_{2})}$ independent codewords $V^{n}$ according to $\prod_{t=1}^{n}P(v_{t}|w_{t}(m_{00}))$. Label these codewords $V^{n}(m_{00},m_{22},b_{2})$ where $m_{22}\in[1:2^{nR_{22}}]$ and $b_{2}\in[1:2^{nB_{2}}]$.
  4. 4.
    Given a triple $(m_{00},m_{11},m_{22})$ to be transmitted, find a pair of indices $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ that satisfies
    $(W^{n}(m_{00}),U^{n}(m_{00},m_{11},b_{1}^{\mathcal{T}}),V^{n}(m_{00},m_{22},b_{2}^{\mathcal{T}}))\\ \qquad\in\mathcal{T}_{\epsilon}^{n}(P_{WUV}).$ (172)
    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (172), let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}+b_{2}$. If there is no such a pair, declare an encoding error. Based on the mutual covering lemma [29], one can easily see that for sufficiently large $n$ and small $\epsilon$, the probability of encoding error tends to zero provided that
    $B_{1}+B_{2}\geq I(U;V|W).$ (173)
    The transmitter then generates a codeword $X^{n}$ according to
    $\prod_{t=1}^{n}P(x_{t}|w_{t}(m_{00}),u_{t}(m_{00},m_{11},b_{1}^{\mathcal{T}}),v_{t}(m_{00},m_{22},b_{2}^{\mathcal{T}}))$
    and sends it over the channel.
B-B Encoding and Decoding at the Receivers
  1. 1.
    Generate at random $2^{n\hat{R}_{2}}$ independent codewords $\hat{Y}_{2}^{n}$ according to the distribution $\prod_{t=1}^{n}P(\hat{y}_{2,t})$. Label these codewords $\hat{Y}_{2}^{n}(\hat{m}_{2})$ where $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$.
  2. 2.
    Partition the set $[1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21})}$ elements. For a given $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$, let $\mathfrak{C}_{21}(\hat{m}_{2})$ denote the index of the bin that $\hat{m}_{2}$ belongs to.
  3. 3.
    Given the output sequence $Y_{2}^{n}$ at the second receiver, let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies
    $(Y_{2}^{n},\hat{Y}_{2}^{n}(\hat{m}_{2}^{\mathcal{T}}))\in\mathcal{T}_{\epsilon}^{n}(P_{Y_{2}\hat{Y}_{2}}).$ (174)
    If there is no integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies (174), then declare an error. The probability of this error tends to zero provided that
    $\hat{R}_{2}\geq I(Y_{2};\hat{Y}_{2}).$ (175)
    The receiver $Y_{2}$ then sends $\kappa=\mathfrak{C}_{21}(\hat{m}_{2}^{\mathcal{T}})$ to the receiver $Y_{1}$ through the digital link.
  4. 4.
    Given its output sequence $Y_{1}^{n}$ and $\kappa\in[1:2^{n\hat{C}_{21}}]$, the first receiver seeks a unique triple $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ for which there exists some $\hat{Y}_{2}^{n}$ that satisfies
    $(W^{n}(m_{00}^{*}),U^{n}(m_{00}^{*},m_{11}^{*},b_{1}^{*}),Y_{1}^{n},\hat{Y}_{2}^{n})\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WUY_{1}\hat{Y}_{2}})$ (176)
    and
    $\mathfrak{C}_{21}(\hat{Y}_{2}^{n})=\kappa.$ (177)
    If there is such a unique triple, the first receiver decodes its intended messages as $m_{00}^{*}$ and $m_{11}^{*}$. Otherwise, it declares an error. The probability that the actual transmitted messages do not satisfy (176)-(177) is negligible for sufficiently large $n$. Consider the error event that some messages other than the actual transmitted ones together with some $\hat{Y}_{2}^{n}$ would satisfy (176)-(177). This event comprises two cases. First, this $\hat{Y}_{2}^{n}$ is not the codeword actually selected from the quantization codebook, i.e., $\hat{Y}_{2}^{n}(\hat{m}_{2}^{\mathcal{T}})$. Second, this $\hat{Y}_{2}^{n}$ is actually $\hat{Y}_{2}^{n}(\hat{m}_{2}^{\mathcal{T}})$. Using the multivariate packing lemma [29], one can show that the probability of the first case tends to zero provided that
    $\hat{R}_{2}-\hat{C}_{21}+$ $R_{11}+B_{1}\leq$
    $I(U;Y_{1}|W)+I(W,U,Y_{1};\hat{Y}_{2})$ (178)
    $\hat{R}_{2}-\hat{C}_{21}+$ $R_{00}+R_{11}+B_{1}\leq$
    $I(W,U;Y_{1})+I(W,U,Y_{1};\hat{Y}_{2})$ (179)
    and the probability of the second case tends to zero provided that
    $R_{11}+B_{1}\leq I(U;Y_{1},\hat{Y}_{2}|W)$ (180)
    $R_{00}+R_{11}+B_{1}\leq I(W,U;Y_{1},\hat{Y}_{2}).$ (181)
  5. 5.
    Generate at random $2^{n\hat{R}_{1}}$ independent codewords $\hat{Y}_{1}^{n}$ according to the distribution $\prod_{t=1}^{n}P(\hat{y}_{1,t})$. Label these codewords $\hat{Y}_{1}^{n}(\hat{m}_{1})$ where $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$.
  6. 6.
    Partition the set $[1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}^{q}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12}^{q})}$ elements. For a given $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$, let $\mathfrak{C}_{12}^{q}(\hat{m}_{1})$ denote the index of the bin that $\hat{m}_{1}$ belongs to.
  7. 7.
    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies
    $(W^{n}(m_{00}^{*}),U^{n}(m_{00}^{*},m_{11}^{*},b_{1}^{*}),Y_{1}^{n},\hat{Y}_{1}^{n}(\hat{m}_{1}^{\mathcal{T}}))\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WUY_{1}\hat{Y}_{1}}).$ (182)
    Note that $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ is the unique triple decoded in Step 4 according to (176)-(177). If there is no integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies (182), then declare an error. The probability of this error event tends to zero provided that
    $\hat{R}_{1}\geq I(W,U,Y_{1};\hat{Y}_{1}).$ (183)
  8. 8.
    Partition the set $[1:2^{nR_{00}}]$ into $2^{n\hat{C}_{12}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{12}^{d})}$ elements. For a given $m_{00}\in[1:2^{nR_{00}}]$, let $\mathfrak{C}_{12}^{d}(m_{00})$ denote the index of the bin that $m_{00}$ belongs to.
  9. 9.
    The receiver $Y_{1}$ then sends the pair $(\theta_{1},\theta_{2})=(\mathfrak{C}_{12}^{d}(m_{00}^{*}),\mathfrak{C}_{12}^{q}(\hat{m}_{1}^{\mathcal{T}}))$ to the receiver $Y_{2}$ through the digital link. (Note that $m_{00}^{*}$ is the decoded message in Step 4 at the receiver $Y_{1}$.)
  10. 10.
    Given $(\theta_{1},\theta_{2})\in[1:2^{n\hat{C}_{12}^{d}}]\times[1:2^{\hat{C}_{12}^{q}}]$ and its output sequence $Y_{2}^{n}$, the second receiver seeks a unique triple $(m_{00}^{\bullet},m_{22}^{\bullet},b_{2}^{\bullet})$ for which there exists some $\hat{Y}_{1}^{n}$ that satisfies
    $(W^{n}(m_{00}^{\bullet}),V^{n}(m_{00}^{\bullet},m_{22}^{\bullet},b_{2}^{\bullet}),\hat{Y}_{1}^{n},Y_{2}^{n})\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WV\hat{Y}_{1}Y_{2}}),$ (184)
    and
    $\mathfrak{C}_{12}^{d}(m_{00}^{\bullet})$ $=\theta_{1},$ (185)
    $\mathfrak{C}_{12}^{q}(\hat{Y}_{1}^{n})$ $=\theta_{2}.$ (186)
    If there is such a unique triple, the second receiver decodes the its intended messages as $m_{00}^{\bullet}$ and $m_{22}^{\bullet}$. Otherwise, it declares an error. The probability that the actual transmitted messages do not satisfy (184)-(186) is negligible for sufficiently large $n$. Consider the error event that some messages other than the actual transmitted ones together with some $\hat{Y}_{1}^{n}$ would satisfy (184)-(186). This event comprises two cases. First, this $\hat{Y}_{1}^{n}$ is not the codeword actually selected from the quantization codebook, i.e., $\hat{Y}_{1}^{n}(\hat{m}_{1}^{\mathcal{T}})$. Second, this $\hat{Y}_{1}^{n}$ is actually $\hat{Y}_{1}^{n}(\hat{m}_{1}^{\mathcal{T}})$. Using the multivariate packing lemma [29], one can show that the probability of the first case tends to zero provided that
    $\hat{R}_{1}-\hat{C}_{12}^{q}+$ $R_{22}+B_{2}\leq$
    $I(V;Y_{2}|W)+I(W,V,Y_{2};\hat{Y}_{1})$ (187)
    $\hat{R}_{1}-\hat{C}_{12}^{q}+$ $R_{00}-\hat{C}_{12}^{d}+R_{22}+B_{2}\leq$
    $I(W,V;Y_{2})+I(W,V,Y_{2};\hat{Y}_{1}),$ (188)
    and the probability of the second case tends to zero provided that
    $R_{22}+B_{2}\leq I(V;\hat{Y}_{1},Y_{2}|W)$ (189)
    $R_{00}-\hat{C}_{12}^{d}+R_{22}+B_{2}\leq I(W,V;\hat{Y}_{1},Y_{2}).$ (190)
B-C Final Analysis
Using the Fourier-Motzkin algorithm to eliminate $\hat{R}_{2}$ and $\hat{C}_{21}$ and by algebraic computation, one can see that (171), (175), (4)-(179), and (180)-(181) can be equivalently rewritten in the following form:
$R_{11}+B_{1}$ $\leq I(U;Y_{1}|W)+I(W,U,Y_{1};\hat{Y}_{2})$
$\qquad\qquad+C_{21}-I(Y_{2};\hat{Y}_{2})$ (191)
$R_{00}+R_{11}+B_{1}$ $\leq I(W,U;Y_{1})+I(W,U,Y_{1};\hat{Y}_{2})$
$\qquad\qquad+C_{21}-I(Y_{2};\hat{Y}_{2})$ (192)
$R_{11}+B_{1}$ $\leq I(U;Y_{1},\hat{Y}_{2}|W)$ (193)
$R_{00}+R_{11}+B_{1}$ $\leq I(W,U;Y_{1},\hat{Y}_{2}).$ (194)
Moreover, the Markov chain condition $(U,W,Y_{1})\rightarrow Y_{2}\rightarrow\hat{Y}_{2}$ on the joint distribution implies that
$I(W,U,Y_{1};\hat{Y}_{2})+C_{21}-I(Y_{2};\hat{Y}_{2})=\\ C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}).$ (195)
One can easily see that if $C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$. Otherwise, we keep $C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$ as is. Then, (B-C)-(194) are equivalent to
$R_{11}$ $+B_{1}\leq\min\Big{\{}I(U;Y_{1}|W)$
$+$ $\{C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+},I(U;Y_{1},\hat{Y}_{2}|W)\Big{\}}$ (196)
$R_{00}$ $+R_{11}+B_{1}\leq\min\Big{\{}I(W,U;Y_{1})$
$+$ $\{C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+},I(W,U;Y_{1},\hat{Y}_{2})\Big{\}}.$ (197)
Similarly, one can see that (169)-(170), (183), (10)-(188), and (189)-(190) can be equivalently rewritten as
$R_{22}+$ $B_{2}\leq I(V;Y_{2}|W)+I(W,V,Y_{2};\hat{Y}_{1})$
$+{\alpha}_{1}C_{12}-I(W,U,Y_{1};\hat{Y}_{1})$ (198)
$R_{00}+R_{22}+$ $B_{2}\leq I(W,V;Y_{2})+I(W,V,Y_{2};\hat{Y}_{1})$
$+{\alpha}_{1}C_{12}-I(W,U,Y_{1};\hat{Y}_{1})+\bar{\alpha}_{1}{C}_{12}$ (199)
$R_{22}+$ $B_{2}\leq I(V;\hat{Y}_{1},Y_{2}|W)$ (200)
$R_{00}+R_{22}+$ $B_{2}\leq I(W,V;\hat{Y}_{1},Y_{2})+\bar{\alpha}_{1}{C}_{12}.$ (201)
In addition, the Markov chain condition $(V,Y_{2})\rightarrow(Y_{1},U,W)\rightarrow\hat{Y}_{1}$ on the joint distribution implies that
$I(W,V,Y_{2};\hat{Y}_{1})+{\alpha}_{1}C_{12}-I(W,U,Y_{1};\hat{Y}_{1})=\\ \alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2}).$ (202)
If $\alpha_{1}C_{12}<I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})$, then one can easily show that the optimal $\hat{Y}_{1}=\varnothing$. Therefore, (B-C)-(201) can be given in the following equivalent form:
$R_{22}$ $+B_{2}\leq\min\Big{\{}I(V;Y_{2}|W)+\{\alpha_{1}C_{12}-$
$I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+},I(V;\hat{Y}_{1},Y_{2}|W)\Big{\}}$ (203)
$R_{00}+R_{22}$ $+B_{2}\leq\min\Big{\{}I(W,V;Y_{2})$
+$+\{\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}+\bar{\alpha}_{1}{C}_{12},$
$\ I(W,V;\hat{Y}_{1},Y_{2})+\bar{\alpha}_{1}{C}_{12}\Big{\}}.$ (204)
Lastly, by considering the bounds (173), (B-C)-(197), and (B-C)-(204) and by applying the Fourier–Motzkin algorithm to eliminate all the partial rates, one can derive the rate region $\mathcal{R}_{i}^{(1)}$. The proof is thus complete.
We note that the error analysis here can also be derived from the general inner bounds of [33], but the general expressions of [33] are not always easy to simplify.
Appendix C Proof of Proposition 2
Fix a joint distribution
$P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|w,y_{2})$
and let $\alpha_{2}\in[0,1]$. Similar to the proof of Proposition 1, we construct random codebooks of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively.
Each of the private messages $M_{1}$ and $M_{2}$ is split into two parts as follows:
$M_{1}$ $\longmapsto(M_{10},M_{11}),$ (205)
where $M_{10}\in[1:2^{nR_{10}}]$, $M_{11}\in[1:2^{nR_{11}}]$, and
$M_{2}$ $\longmapsto(M_{20},M_{22}),$ (206)
where $M_{20}\in[1:2^{nR_{20}}]$, $M_{22}\in[1:2^{nR_{22}}]$. Therefore,
$R_{1}=R_{10}+R_{11},$ (207)
$R_{2}=R_{20}+R_{22}.$ (208)
The sub-messages $M_{10}$ and $M_{20}$ are decoded at both receivers, i.e., the triple $(M_{0},M_{10},M_{20})$ is considered as a common message with rate
$R_{00}=R_{0}+R_{10}+R_{20}.$ (209)
Moreover, let $B_{1},B_{2},\hat{R}_{1},\hat{R}_{2}$ be non-negative real numbers denoting the rates of the binning steps and the compression steps in the proof. Also, define
$\hat{C}_{12}$ $=\min\{C_{12},\hat{R}_{1}\}$ (210)
$\hat{C}_{21}^{d}$ $=\min\{\bar{\alpha}_{2}C_{21},R_{00}\}$ (211)
$\hat{C}_{21}^{q}$ $=\min\{{\alpha}_{2}C_{21},\hat{R}_{2}\}$ (212)
as the portion of the conferencing link rates used in the two directions and for decode-and-forward and quantize-bin-and-forward, respectively.
C-A Encoding at the Transmitter
  1. 1.
    Generate at random $2^{nR_{00}}$ independent codewords $W^{n}$ according to $P(w^{n})=\prod_{t=1}^{n}P(w_{t})$. Label these codewords $W^{n}(m_{00})$ where $m_{00}\in[1:2^{nR_{00}}]$.
  2. 2.
    For each codeword $W^{n}(m_{00})$, generate at random $2^{n(R_{11}+B_{1})}$ independent codewords $U^{n}$ according to $\prod_{t=1}^{n}P(u_{t}|w_{t}(m_{00}))$. Label these codewords $U^{n}(m_{00},m_{11},b_{1})$ where $m_{11}\in[1:2^{nR_{11}}]$ and $b_{1}\in[1:2^{nB_{1}}]$.
  3. 3.
    For each codeword $W^{n}(m_{00})$, generate at random $2^{n(R_{22}+B_{2})}$ independent codewords $V^{n}$ according to $\prod_{t=1}^{n}P(v_{t}|w_{t}(m_{00}))$. Label these codewords $V^{n}(m_{00},m_{22},b_{2})$ where $m_{22}\in[1:2^{nR_{22}}]$ and $b_{2}\in[1:2^{nB_{2}}]$.
  4. 4.
    Given a triple $(m_{00},m_{11},m_{22})$ to be transmitted, let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be a pair that satisfies
    $(W^{n}(m_{00}),U^{n}(m_{00},m_{11},b_{1}^{\mathcal{T}}),V^{n}(m_{00},m_{22},b_{2}^{\mathcal{T}}))\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WUV}).$ (213)
    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (213), then let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}+b_{2}$. If there is no such a pair, declare encoding error. Based on the mutual covering lemma [29], one can easily see that for sufficiently large $n$ and small $\epsilon$, the probability of encoding error tends to zero provided that
    $B_{1}+B_{2}\geq I(U;V|W).$ (214)
    The transmitter then generates a codeword $X^{n}$ according to
    $\prod_{t=1}^{n}P(x_{t}|w_{t}(m_{00}),u_{t}(m_{00},m_{11},b_{1}^{\mathcal{T}}),v_{t}(m_{00},m_{22},b_{2}^{\mathcal{T}}))$
    and sends it over the channel.
C-B Encoding and Decoding at the Receivers
  1. 1.
    Partition the set $[1:2^{nR_{00}}]$ into $2^{n\hat{C}_{21}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{21}^{d})}$ elements. For a given $m_{00}\in[1:2^{nR_{00}}]$, let $\mathfrak{C}_{21}^{d}(m_{00})$ denote the index of the bin that $m_{00}$ belongs to.
  2. 2.
    Generate at random $2^{n\hat{R}_{2}}$ independent codewords $\hat{Y}_{2}^{n}$ according to the distribution $\prod_{t=1}^{n}P(\hat{y}_{2,t})$. Label these codewords $\hat{Y}_{2}^{n}(\hat{m}_{2})$ where $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$.
  3. 3.
    Partition the set $[1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}^{q}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21}^{q})}$ elements. For a given $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$, let $\mathfrak{C}_{21}^{q}(\hat{m}_{2})$ denote the index of the bin that $\hat{m}_{2}$ belongs to.
  4. 4.
    Given its output sequence $Y_{2}^{n}$, the second receiver seeks a unique $m_{00}^{\bullet}$ for which there exists some $(m_{22},b_{2})$ that satisfies
    $(W^{n}(m_{00}^{\bullet}),V^{n}(m_{00}^{\bullet},m_{22},b_{2}),Y_{2}^{n})\in\mathcal{T}_{\epsilon}^{n}(P_{WVY_{2}}).$ (215)
    If there is no such $m_{00}^{\bullet}$ or if there is more than one, declare an error. The probability of this error event tends to zero provided that
    $R_{00}+R_{22}+B_{2}\leq I(W,V;Y_{2})$ (216)
  5. 5.
    Let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies
    $(W^{n}(m_{00}^{\bullet}),Y_{2}^{n},\hat{Y}_{2}^{n}(\hat{m}_{2}^{\mathcal{T}}))\in\mathcal{T}_{\epsilon}^{n}(P_{WY_{2}\hat{Y}_{2}}).$ (217)
    If there is no integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies (217), then declare an error. The probability of this error tends to zero provided that
    $\hat{R}_{2}\geq I(W,Y_{2};\hat{Y}_{2})$ (218)
    The receiver $Y_{2}$ then sends the pair $(\theta_{1},\theta_{2})=(\mathfrak{C}_{21}^{d}(m_{00}^{\bullet}),\mathfrak{C}_{21}^{q}(\hat{m}_{2}^{\mathcal{T}}))$ to the receiver $Y_{1}$ through the digital link.
  6. 6.
    Given $(\theta_{1},\theta_{2})\in[1:2^{n\hat{C}_{21}^{d}}]\times[1:2^{n\hat{C}_{21}^{q}}]$ and its output sequence $Y_{1}^{n}$, the first receiver seeks a unique triple $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ for which there exists some $\hat{Y}_{2}^{n}$ that satisfies
    $(W^{n}(m_{00}^{*}),U^{n}(m_{00}^{*},m_{11}^{*},b_{1}^{*}),Y_{1}^{n},\hat{Y}_{2}^{n})\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WUY_{1}\hat{Y}_{2}})$ (219)
    and
    $\mathfrak{C}_{21}^{d}(m_{00}^{*})$ $=\theta_{1},$ (220)
    $\mathfrak{C}_{21}^{q}(\hat{Y}_{2}^{n})$ $=\theta_{2}.$ (221)
    If there is such a unique triple, the first receiver decodes its intended messages as $m_{00}^{*}$ and $m_{11}^{*}$. Otherwise, it declares an error. The probability that the actual transmitted messages do not satisfy (219)-(221) is negligible for sufficiently large $n$. Consider the error event that some messages other than the actual transmitted ones together with some $\hat{Y}_{2}^{n}$ would satisfy (219)-(221). This event comprises two cases. First, this $\hat{Y}_{2}^{n}$ is not the codeword actually selected from the quantization codebook, i.e., $\hat{Y}_{2}^{n}(\hat{m}_{2}^{\mathcal{T}})$. Second, this $\hat{Y}_{2}^{n}$ is actually $\hat{Y}_{2}^{n}(\hat{m}_{2}^{\mathcal{T}})$. Using the multivariate packing lemma [29], one can see that the probability of the first case tends to zero provided that
    $\hat{R}_{2}-\hat{C}_{21}^{q}+$ $R_{11}+B_{1}\leq$
    $I(U;Y_{1}|W)+I(W,U,Y_{1};\hat{Y}_{2})$ (222)
    $\hat{R}_{2}-\hat{C}_{21}^{q}+$ $R_{00}-\hat{C}_{21}^{d}+R_{11}+B_{1}\leq$
    $I(W,U;Y_{1})+I(W,U,Y_{1};\hat{Y}_{2})$ (223)
    and the probability of the second case tends to zero provided that
    $R_{11}+B_{1}\leq I(U;Y_{1},\hat{Y}_{2}|W)$ (224)
    $R_{00}-\hat{C}_{21}^{d}+R_{11}+B_{1}\leq I(W,U;Y_{1},\hat{Y}_{2}).$ (225)
  7. 7.
    Generate at random $2^{n\hat{R}_{1}}$ independent codewords $\hat{Y}_{1}^{n}$ according to the distribution $\prod_{t=1}^{n}P(\hat{y}_{1,t})$. Label these codewords $\hat{Y}_{1}^{n}(\hat{m}_{1})$ where $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$.
  8. 8.
    Partition the set $[1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12})}$ elements. For a given $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$, let $\mathfrak{C}_{12}(\hat{m}_{1})$ denote the index of the bin that $\hat{m}_{1}$ belongs to.
  9. 9.
    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies
    $(W^{n}(m_{00}^{*}),U^{n}(m_{00}^{*},m_{11}^{*},b_{1}^{*}),Y_{1}^{n},\hat{Y}_{1}^{n}(\hat{m}_{1}^{\mathcal{T}}))\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WUY_{1}\hat{Y}_{1}}),$ (226)
    where $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ is the unique triple decoded in Step 6 according to (219)-(221). If there is no integer $[1:2^{n\hat{R}_{1}}]$ that satisfies (226), then declare an error. The probability of this error event tends to zero provided that
    $\hat{R}_{1}\geq I(W,U,Y_{1};\hat{Y}_{1})$ (227)
    The receiver $Y_{1}$ then sends $\kappa=\mathfrak{C}_{12}(\hat{m}_{1}^{\mathcal{T}})$ to the receiver $Y_{2}$ through the digital link.
  10. 10.
    Given $\kappa\in[1:2^{n\hat{C}_{12}}]$ and its output sequence $Y_{2}^{n}$, the second receiver seeks a unique pair $(m_{22}^{\bullet},b_{2}^{\bullet})$ for which there exists some $\hat{Y}_{1}^{n}$ that satisfies
    $(W^{n}(m_{00}^{\bullet}),V^{n}(m_{00}^{\bullet},m_{22}^{\bullet},b_{2}^{\bullet}),\hat{Y}_{1}^{n},Y_{2}^{n})\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WV\hat{Y}_{1}Y_{2}})$ (228)
    and
    $\mathfrak{C}_{12}(\hat{Y}_{1}^{n})=\kappa.$ (229)
    Note that $m_{00}^{\bullet}$ has been already decoded in Step 4. If there is such a unique pair, the receiver decodes its intended (private) message as $m_{22}^{\bullet}$. Otherwise, it declares an error. The probability that the actual transmitted messages do not satisfy (228)-(229) is negligible for sufficiently large $n$. Consider the error event that some messages other than the actual transmitted ones together with some $\hat{Y}_{1}^{n}$ would satisfy (228)-(229). This event comprises two cases. First, this $\hat{Y}_{1}^{n}$ is not the codeword actually selected from the quantization codebook, i.e., $\hat{Y}_{1}^{n}(\hat{m}_{1}^{\mathcal{T}})$. Second, this $\hat{Y}_{1}^{n}$ is actually $\hat{Y}_{1}^{n}(\hat{m}_{1}^{\mathcal{T}})$. Using the multivariate packing lemma [29], one can show that the probability of the first case tends to zero provided that
    $\hat{R}_{1}-\hat{C}_{12}+R_{22}+B_{2}\leq I(V;Y_{2}|W)+I(W,V,Y_{2};\hat{Y}_{1})$ (230)
    and the probability of the second case tends to zero provided that
    $R_{22}+B_{2}\leq I(V;\hat{Y}_{1},Y_{2}|W).$ (231)
C-C Final Analysis
By using Fourier-Motzkin elimination and by algebraic computation, one can see that (210)-(212), (218), and (222)-(225) can be equivalently rewritten in the following form:
$R_{11}+B_{1}$ $\leq I(U;Y_{1}|W)+I(W,U,Y_{1};\hat{Y}_{2})$
$\ \ +{\alpha}_{2}C_{21}-I(W,Y_{2};\hat{Y}_{2})$ (232)
$R_{00}+R_{11}+B_{1}$ $\leq I(W,U;Y_{1})+I(W,U,Y_{1};\hat{Y}_{2})$
$\ \ +{\alpha}_{2}C_{21}-I(W,Y_{2};\hat{Y}_{2})+\bar{\alpha}_{2}C_{21}$ (233)
$R_{11}+B_{1}$ $\leq I(U;Y_{1},\hat{Y}_{2}|W)$ (234)
$R_{00}+R_{11}+B_{1}$ $\leq I(W,U;Y_{1},\hat{Y}_{2})+\bar{\alpha}_{2}C_{21}.$ (235)
Moreover, the Markov chain condition $(U,Y_{1})\rightarrow(W,Y_{2})\rightarrow\hat{Y}_{2}$ on the joint distribution implies that
$I(W,U,Y_{1};\hat{Y}_{2})+{\alpha}_{2}C_{21}-I(W,Y_{2};\hat{Y}_{2})=\\ {\alpha}_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}).$ (236)
One can easily see that if ${\alpha}_{2}C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$. Thus, (232)-(235) are equivalent to
$R_{11}+B_{1}$ $\leq\min\Big{\{}I(U;Y_{1}|W)$
+$\qquad+\{{\alpha}_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+},$
$\qquad\qquad I(U;Y_{1},\hat{Y}_{2}|W)\Big{\}}$ (237)
$R_{00}+R_{11}+B_{1}$ $\leq\min\Big{\{}I(W,U;Y_{1})+\{{\alpha}_{2}C_{21}$
$\qquad-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}+\bar{\alpha}_{2}C_{21},$
$\qquad\qquad I(W,U;Y_{1},\hat{Y}_{2})+\bar{\alpha}_{2}C_{21}\Big{\}}.$ (238)
Similarly, by algebraic computation, one can see that (210)-(212), (227), (230), and (231) can be rewritten in the following equivalent form:
$R_{22}+B_{2}$ $\leq I(V;Y_{2}|W)+I(W,V,Y_{2};\hat{Y}_{1})+C_{12}$
$\qquad\qquad-I(W,U,Y_{1};\hat{Y}_{1})$ (239)
$R_{22}+B_{2}$ $\leq I(V;\hat{Y}_{1},Y_{2}|W)$ (240)
$R_{00}+R_{22}+B_{2}$ $\leq I(W,V;Y_{2}).$ (241)
The Markov chain condition $(V,Y_{2})\rightarrow(U,W,Y_{1})\rightarrow\hat{Y}_{1}$ on the joint distribution implies that
$I(W,V,Y_{2};\hat{Y}_{1})+C_{12}-I(W,U,Y_{1};\hat{Y}_{1})=\\ C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2}).$ (242)
If $C_{12}<I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})$, then the optimal $\hat{Y}_{1}=\varnothing$. Therefore, (239)-(241) are equivalent to
$R_{22}+B_{2}$ $\leq\min\Big{\{}I(V;Y_{2}|W)$
+$\qquad+\{C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+},$
$\quad\qquad I(V;\hat{Y}_{1},Y_{2}|W)\Big{\}}$ (243)
$R_{00}+R_{22}+B_{2}$ $\leq I(W,V;Y_{2}).$ (244)
Finally, by considering the bounds (237)-(238) and (243)-(244) and by applying the Fourier–Motzkin elimination algorithm, we obtain the rate region $\mathcal{R}_{i}^{(2)}$. The proof is thus complete.
Appendix D Proof of Theorem 6
We present the optimization over the auxiliary random variable $V$ in $\mathcal{R}_{o}$ in Theorem 1. The constraints involving the auxiliary random variable $U$ can be evaluated similarly. First note that since $|a|\geq|b|$, the receiver $Y_{1}$ is less noisy than the receiver $Y_{2}$, i.e.,
$I(V;Y_{2})\leq I(V;Y_{1}).$ (245)
Therefore, we have
$I($ $X;Y_{1}|Y_{2},V)+I(X;Y_{2})$
$=I(X;Y_{1}|Y_{2},V)+I(X;Y_{2}|V)+I(V;Y_{2})$ (246)
$=I(X;Y_{1},Y_{2}|V)+I(V;Y_{2})$ (247)
$\leq I(X;Y_{1},Y_{2}|V)+I(V;Y_{1})$ (248)
$=I(X;Y_{2}|Y_{1},V)+I(X;Y_{1}|V)+I(V;Y_{1})$ (249)
$=I(X;Y_{2}|Y_{1},V)+I(X;Y_{1}).$ (250)
In other words, constraint (3) is inactive. Considering (2), (4), (7), (9), and (11) and accounting for the Markov chain $V\rightarrow X\rightarrow(Y_{1},Y_{2})$, it is easy to see that we need to maximize the following terms simultaneously:
$I_{1}$ $\doteq I(X;Y_{1},Y_{2}|V)+I(V;Y_{2})$ (251)
$I_{2}$ $\doteq I(V;Y_{2})$ (252)
$I_{3}$ $\doteq I(X;Y_{1}|V)+I(V;Y_{2})$ (253)
$I_{4}$ $\doteq I(X;Y_{1},Y_{2}).$ (254)
Consider first the special cases where $\lambda=\frac{b}{a}$. In this case, $Z_{1}$ and $Z_{2}$ can be thought of as being related as
$Z_{2}=\frac{b}{a}Z_{1}+\tilde{Z},$ (255)
where $Z_{1}$, $\tilde{Z}$ are independent. Since $Y_{1}=aX+Z_{1}$, we have
$Y_{2}$ $=bX+Z_{2}$ (256)
$=bX+\frac{b}{a}Z_{1}+\tilde{Z}$ (257)
$=\frac{b}{a}Y_{1}+\tilde{Z}.$ (258)
So, the BC is physically degraded, i.e., $X\rightarrow Y_{1}\rightarrow Y_{2}$ forms a Markov chain. The simultaneous maximization of the four mutual information terms (251)-(254) then reduces to the simultaneous maximization of $h(Y_{1}|V)$ and $h(Y_{1})$ together with minimization of $h(Y_{2}|V)$, which is achieved by jointly Gaussian $(X,V)$ as can be shown by the entropy power inequality for the degraded Gaussian broadcast channel [29].
In the rest of the proof, we assume $\lambda\neq\frac{b}{a}$. Define a new virtual output $\hat{Y}$ (which is a scaled version of the minimum mean-squared error (MMSE) estimate of $X$ given $(Y_{1},Y_{2})$) as follows:
$\hat{Y}$ $\doteq(a-\lambda b)Y_{1}+(b-\lambda a)Y_{2}$ (259)
$=\left((a-\lambda b)a+(b-\lambda a)b\right)X+\hat{Z}$ (260)
$=(a^{2}+b^{2}-2\lambda ab)X+\hat{Z}$ (261)
where
$\hat{Z}\doteq(a-\lambda b)Z_{1}+(b-\lambda a)Z_{2}.$ (262)
When $\lambda\neq\frac{b}{a}$, it is clear that there is a one-to-one mapping between $(\hat{Y},Y_{2})$ and $(Y_{1},Y_{2})$. Moreover, $\lambda\neq\frac{b}{a}$ and the fact $|\lambda|\leq 1$ imply that $a^{2}+b^{2}-2\lambda ab\neq 0$, so we can write
$Y_{2}=\frac{b}{a^{2}+b^{2}-2\lambda ab}\hat{Y}+\hat{Z_{2}},$ (263)
where
$\hat{Z_{2}}\doteq\frac{-b\hat{Z}}{a^{2}+b^{2}-2\lambda ab}+Z_{2}.$ (264)
It can be verified that $\hat{Z_{2}}$ and $\hat{Z}$ are uncorrelated, and therefore they are independent (as they are jointly Gaussian). This means that the output $Y_{2}$ is in fact a noisy version of $\hat{Y}$, i.e., $X\rightarrow\hat{Y}\rightarrow Y_{2}$ forms a Markov chain. This relationship can also be recognized from the fact that $\hat{Y}$ is a (scaled) MMSE estimate of $X$ given $(Y_{1},Y_{2})$. Thus, we have
$I(X;Y_{1},Y_{2}|V)$ $=I(X;\hat{Y},Y_{2}|V)=I(X;\hat{Y}|V),$ (265)
$I(X;Y_{1},Y_{2})$ $=I(X;\hat{Y},Y_{2})=I(X;\hat{Y}).$ (266)
Now, utilizing (265)-(266) for bounding the maximum values of (251)-(254), we see that after expressing the mutual information in terms of differential entropy, the optimization is effectively to maximize $h(\hat{Y}|V)$, $h(Y_{1}|V)$ and $h(\hat{Y})$, and simultaneously to minimize $h(Y_{2}|V)$.
We complete this step using the entropy power inequality. First,
$\frac{1}{2}\log(2\pi e)$ $=h(Z_{1})$ (267)
$=h(Y_{1}|X,V)$ (268)
$\leq h(Y_{1}|V)$ (269)
$\leq h(Y_{1})$ (270)
$\leq\frac{1}{2}\log 2\pi e(a^{2}P+1).$ (271)
By comparing the two sides of (269)-(271), we find that there is $\beta\in[0,1]$ such that
$h(Y_{1}|V)=\frac{1}{2}\log 2\pi e(\beta a^{2}P+1).$ (272)
Now, let $\hat{Z}^{*}$ be a Gaussian random variable independent of all other variables and with zero mean and a variance equal to $\frac{(a\lambda-b)^{2}}{\mu}$ where $\mu\doteq a^{2}+b^{2}-2\lambda ab$. We then have
$h(\hat{Y}|V)$ $=h(\mu X+\hat{Z})$ (273)
$=h\left(aX+\frac{a}{\mu}\hat{Z}\right)-\frac{1}{2}\log\left(\frac{a^{2}}{\mu^{2}}\right)$ (274)
$\leq\frac{1}{2}\log\left(e^{2h(aX+\frac{a}{\mu}\hat{Z}+\hat{Z}^{*}|V)}-e^{2h(\hat{Z}^{*}|V)}\right)$
$\qquad\qquad\qquad-\frac{1}{2}\log\left(\frac{a^{2}}{\mu^{2}}\right)$ (275)
$=\frac{1}{2}\log\left(e^{2h(aX+Z_{1}|V)}-e^{2h(\hat{Z}^{*})}\right)$
$\qquad\qquad\qquad-\frac{1}{2}\log\left(\frac{a^{2}}{\mu^{2}}\right)$ (276)
$=\frac{1}{2}\log\left(e^{2h(Y_{1}|V)}-e^{2h(\hat{Z}^{*})}\right)$
$\qquad\qquad\qquad-\frac{1}{2}\log\left(\frac{a^{2}}{\mu^{2}}\right)$ (277)
$=\frac{1}{2}\log\left(2\pi e\left(\beta\mu P+1-\lambda^{2}\right)\mu\right)$ (278)
where (275) is due to the entropy power inequality, and (276) holds because $\frac{a}{\mu}\hat{Z}+\hat{Z}^{*}$ is a Gaussian random variable with zero mean and unit variance and independent of $X$ and $V$, therefore it can be replaced by $Z_{1}$ in the differential entropy expression, and also $\hat{Z}^{*}$ is independent of $V$.
Finally, let $Z_{2}^{*}$ be a Gaussian random variable independent of all other variables and with zero mean and variance $\frac{a^{2}-b^{2}}{b^{2}}$, we have
$h(Y_{2}|V)$ $=h(bX+Z_{2}|V)$
$=h\left(aX+\frac{a}{b}Z_{2}|V\right)-\frac{1}{2}\log\left(\frac{a^{2}}{b^{2}}\right)$ (279)
$=h(aX+Z_{2}+Z_{2}^{*}|V)-\frac{1}{2}\log\left(\frac{a^{2}}{b^{2}}\right)$ (280)
$\geq\frac{1}{2}\log\left(e^{2h(aX+Z_{2}|V)}+e^{2h(Z_{2}^{*}|V)}\right)$
$\qquad\qquad\qquad\qquad-\frac{1}{2}\log\left(\frac{a^{2}}{b^{2}}\right)$ (281)
$=\frac{1}{2}\log\left(e^{2h(aX+Z_{1}|V)}+e^{2h(Z_{2}^{*})}\right)$
$\qquad\qquad\qquad\qquad-\frac{1}{2}\log\left(\frac{a^{2}}{b^{2}}\right)$ (282)
$=\frac{1}{2}\log\left(e^{2h(Y_{1}|V)}+e^{2h(Z_{2}^{*})}\right)$
$\qquad\qquad\qquad\qquad-\frac{1}{2}\log\left(\frac{a^{2}}{b^{2}}\right)$ (283)
$=\frac{1}{2}\log\left(2\pi e\left(\beta b^{2}P+1\right)\right)$ (284)
where (281) is due to the entropy power inequality, and (282) holds because $Z_{2}$ has the same distribution as $Z_{1}$ and both are independent of $(X,V)$, and also $Z_{2}^{*}$ is independent of $V$.
Considering (272), (278), and (284) together with the fact that a Gaussian $X\sim\mathcal{N}(0,P)$ maximizes $h(\hat{Y})$ allows us to conclude that a jointly Gaussian $(X,V)$ simultaneously maximizes the mutual information terms in (251)-(254) and (265)-(266).
The proof can be completed by substituting the evaluation of these mutual information terms (251)-(254) for Gaussian $(X,V)$ into $\mathcal{R}_{o}$ in Theorem 1.

### v4

@@
 Finally, for the general Gaussian BC, a one-sided cooperation scheme based on decode-and-forward from the stronger receiver to the weaker receiver is shown to achieve the capacity region to within $\frac{1}{2}\log(\frac{2}{1-|\lambda|})$ bits, where $\lambda$ is the noise correlation.
-An interesting implication of these results is that for a Gaussian BC with perfectly negatively correlated noises and conferencing decoders with finite cooperation link capacities, it is possible to achieve a strictly positive rate using only infinitesimal amount of transmit power.
+An interesting implication of these results is that for a Gaussian BC with perfectly negatively correlated noises and conferencing decoders with finite cooperation link capacities, it is possible to achieve a strictly positive rate using only an infinitesimal amount of transmit power.
 Index Terms:
-Broadcast channel, bidirectional cooperation, conferencing decoders, capacity region, deterministic channel, Gaussian broadcast channel
+Broadcast channel, capacity region, correlated noise, primitive relay channel, receiver cooperation
 I Introduction
@@
 Consider the case where $Y_{1}=X+Z$ and $Y_{2}=X-Z$, and the conferencing links have infinite capacities $C_{12}=C_{21}=\infty$.
-Then, the BC with conferencing decoders would have infinite capacity at any positive transmit power $P$, because having access to both $Y_{1}$ and $Y_{2}$ allows the noises to be cancelled completely.
+Such a BC with full receiver cooperation has infinite capacity at any positive transmit power $P$, because having access to both $Y_{1}$ and $Y_{2}$ allows the noises to be cancelled completely.
 What happens if the noises are perfectly negatively correlated, but the conferencing links have finite capacity, e.g., $C_{12}=C_{21}=1$ bit per channel use?
-The analysis becomes more delicate and this is one of the main topics of this paper.
-It turns out that an interesting phenomenon emerges—it is possible to achieve a strictly positive rate (i.e., 1 bit in the example above), even with an infinitesimally small transmit power.
+Complete noise cancellation is no longer possible.
+The analysis becomes more delicate, and this is one of the main topics of this paper.
+It turns out that an interesting phenomenon emerges—it is possible to achieve a strictly positive rate (i.e., 1 bit in the example above), even with an infinitesimally small amount of transmit power.
 In contrast to the single-user Gaussian channel where $P\to 0$ implies $C\to 0$, for this Gaussian BC with perfectly negatively correlated noises and conferencing decoders of fixed cooperative link capacities, we can have $P\to 0$, while $C\to 1$.
@@
 Figure 1: The capacity of a Gaussian channel (left) goes to zero if the transmit power $P$ goes to zero at a fixed noise variance $N$.
-In a Gaussian BC with perfectly negatively correlated noises at the receiver (right), conferencing links of fixed capacities $C_{12}$ and $C_{21}$ between the receivers can allow a strictly positive rate to be achieved even at an infinitesimally small transmit power.
+In a Gaussian BC with perfectly negatively correlated noises at the receiver (right), conferencing links of fixed capacities $C_{12}$ and $C_{21}$ between the receivers can allow a strictly positive rate to be achieved even at an infinitesimally small amount of transmit power.
 The above example motivates us to study the BC with conferencing decoders with a specific focus on the Gaussian channel with correlated noises.
@@
 Another achievable region is given in [21] based on coding strategies for the partially cooperative relay broadcast channels.
-A third achievable region is given in [23, Appendix B] for the BC with one-sided cooperation between receivers, which is derived by applying Marton’s coding at the transmitter and decode-and-forward as the cooperative protocol.
+A third achievable region is given in [23, Appendix B] for the BC with one-sided cooperation between the receivers, which is derived by applying Marton’s coding at the transmitter and decode-and-forward as the cooperative protocol.
 All of these achievable schemes are however insufficient to either derive new capacity results for the discrete memoryless BCs, or approximate capacity results for the Gaussian BC.
@@
 For the BC with conferencing decoders, a length-$n$ code with $\Gamma$ conferencing rounds is described as follows.
-The transmitter encodes independent messages $M_{0}$, $M_{1}$, and $M_{2}$, which are uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively, into a codeword and sends over the channel according to the following:
+The transmitter encodes independent messages $M_{0}$, $M_{1}$, and $M_{2}$, which are uniformly distributed over the sets $1:2^{nR_{0}}]$, $1:2^{nR_{1}}]$, and $1:2^{nR_{2}}]$, respectively, into a codeword and sends over the channel according to the following:
 $\Delta:[1:2^{nR_{0}}]\times[1:2^{nR_{1}}]\times[1:2^{nR_{2}}]\longmapsto\mathcal{X}^{n}$
@@
 for some joint PDFs $P(u,v,x)$, where $U,V\rightarrow X\rightarrow Y_{1},Y_{2}$ forms a Markov chain and $|\mathcal{U}|\leq|\mathcal{X}|+2$ and $|\mathcal{V}|\leq|\mathcal{X}|+2$.
-The set $\mathcal{R}_{o}$ constitutes an outer bound on the capacity region.
+Then $\mathcal{R}_{o}$ constitutes an outer bound on the capacity region.
 Proof:
@@
 where
-$\zeta_{1}$ =$=\{\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (17)
-$\zeta_{2}$ =$=\{C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (18)
-for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$ and $\alpha_{1}\in[0,1]$. The convex closure of the set $\mathcal{R}_{i}^{(1)}$ is achievable.
+$\zeta_{1}$ $=\{\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (17)
+$\zeta_{2}$ $=\{C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (18)
+for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$ and $\alpha_{1}\in[0,1]$. The convex closure of $\mathcal{R}_{i}^{(1)}$ is achievable.
 Proof:
@@
 Moreover, after some algebraic computation, one can show that $\mathcal{R}_{i}^{(1)}$, restricted over the distributions of the form $P(u,v,w,x)P(\hat{y}_{1}|w,y_{1})P(\hat{y}_{2}|y_{2})$, includes the achievable rate region of [21, Theorem 3] as a subset.
-Therefore, $\mathcal{R}_{i}^{(1)}$ encompasses these previously known results.
+Therefore, $\mathcal{R}_{i}^{(1)}$ encompasses these previously results.
+The achievable rate region $\mathcal{R}_{i}^{(1)}$ is evaluated over a larger set of distributions, i.e., $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$, as compared to the one in [21, Theorem 3].
+This improvement arises from the novel cooperation scheme at the first receiver, where it transmits a compressed version of the decoded signals U and W along with its received signal $Y_{1}$ to the second receiver.
+As discussed in [30], this novel approach may lead to derivation of new capacity results for certain types of Gaussian channels.
 Let us now specialize the achievable rate region $\mathcal{R}_{i}^{(1)}$ for a primitive relay channel [31].
@@
 $2R_{0}+R_{1}+R_{2}$ $\leq I(U,W;Y_{1})+C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})+I(V,W;Y_{2})-I(U;V|W)$ (47)
-Next, we present a second achievable rate region based on a combination of decode-and-forward and quantize-bin-and-forward at one receiver first and then quantize-bin-and-forward at the other receiver as cooperation strategy.
+Next, we present a second achievable rate region based on a combination of decode-and-forward and quantize-bin-and-forward at one receiver first and then quantize-bin-and-forward at the other receiver as the cooperation strategy.
 This strategy may be advantageous when part of the message is already decodable at the first receiver before conferencing.
@@
 where
-$\eta_{1}$ =$=\{C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (40)
-$\eta_{2}$ =$=\{\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (41)
-for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|w,y_{2})$ and $\alpha_{2}\in[0,1]$. The convex closure of the set $\mathcal{R}_{i}^{(2)}$ is achievable.
+$\eta_{1}$ $=\{C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (40)
+$\eta_{2}$ $=\{\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (41)
+for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|w,y_{2})$ and $\alpha_{2}\in[0,1]$. The convex closure of $\mathcal{R}_{i}^{(2)}$ is achievable.
 Proof:
-The cooperation strategy to achieve the rate region $\mathcal{R}_{i}^{(2)}$ is to apply Marton’s coding with a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{2}$ first and then quantize-bin-and-forward at the receiver $Y_{1}$.
+The cooperation strategy to achieve the rate region $\mathcal{R}_{i}^{(2)}$ is to apply Marton’s coding with a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{2}$ first, then quantize-bin-and-forward at the receiver $Y_{1}$.
 At the receiver $Y_{2}$, the capacity of the conferencing link $C_{21}$ is divided into two parts.
 A fraction $\bar{\alpha}_{2}C_{21}$ is devoted for decode-and-forward and the remaining $\alpha_{2}C_{21}$ is devoted for quantize-bin-and-forward scheme.
-A complete proof of Proposition 2 can be found in Appendix C.
+The details of the proof are in Appendix C.
 ∎
@@
 Remark 5.
-One can obtain an alternative achievable rate region by exchanging the cooperative protocol at the receivers in the acheivability scheme of Proposition 2, i.e., to apply a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{1}$ first and then quantize-bin-and-forward at the receiver $Y_{2}$.
+One can obtain an alternative achievable rate region by exchanging the cooperative protocol at the receivers in the achievability scheme of Proposition 2, i.e., to apply a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{1}$ first and then quantize-bin-and-forward at the receiver $Y_{2}$.
 Similar to the region $\mathcal{R}_{i}^{(1)}$, we can derive the optimal value of the parameter $\alpha_{2}$ for the region $\mathcal{R}_{i}^{(2)}$.
@@
 Proof:
-The achievability is derived by setting $U=X$, $W=V$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$. The converse proof is readily given by $\mathcal{R}_{o}$. ∎
+The achievability is derived by setting $U=X$, $W=V$, $\hat{Y}_{1}=\varnothing$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$. The converse proof is readily given by $\mathcal{R}_{o}$. ∎
 Note that $(Y_{1},Y_{2})$ appears jointly in the first mutual information expression in (60).
@@
 for some $P(v,x)$, where $V\rightarrow X\rightarrow Y_{1},Y_{2}$ forms a Markov chain and $|\mathcal{V}|\leq 3$.
-Figure 3: Capacity region and cut-set outer bound for the BC in Example 2, where $Z$ is a binary noise with ${P(Z=0)=0.2}$.
-Unlike the previous example, where the capacity region coincides with the cut-set bound, the capacity region for this BC is characterized using an auxiliary variable $V$.
+Figure 3: Capacity region and cut-set bound for the BC in Example 2, where $Z$ is a binary noise with P(Z=0)=0.2{P(Z=0)=0.2}.
+Unlike the previous example, for which the capacity region coincides with the cut-set bound, the capacity region for this BC is characterized using an auxiliary variable $V$.
 In Fig.
-3, we plot the capacity region along with the cut-set outer bound for the channel in Example 2 with $P(Z=0)=0.2$ for two cases: $(C_{12},C_{21})=(0,0.9)$ and $(C_{12},C_{21})=(0.1,0.9)$, by searching over all possible $P(v,x)$.
+3, we plot the capacity region along with the cut-set bound for the channel in Example 2 with $P(Z=0)=0.2$ for two cases: $(C_{12},C_{21})=(0,0.9)$ and $(C_{12},C_{21})=(0.1,0.9)$, by searching over all possible $P(v,x)$.
 As can be seen, the capacity region lies strictly below the cut-set bound in this example.
@@
 The optimal cooperation strategy is that the receiver $Y_{2}$ first sends a compressed version of its received signal to the receiver $Y_{1}$ through the digital link $C_{21}$.
-Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal which is in fact the channel noise) and then forwards part of the common message (using the variable $V$) to the receiver $Y_{2}$ through the link $C_{12}$.
-Lastly, the receiver $Y_{2}$ decodes the common message using its received signal.
+Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal which is in fact the channel noise) and then performs decode-and-forward of the common message through the link $C_{12}$ to the receiver $Y_{2}$ (using the variable $V$).
+Lastly, the receiver $Y_{2}$ decodes the common message using also its own signal.
 It is worthwhile to make the following observation.
-For Example 2, as the receiver $Y_{1}$ observes the channel noise only, one might think that a cooperative scheme in which $Y_{1}$ applies quantize-bin-and-forward and $Y_{2}$ applies decode-and-forward is the right strategy.
+In Example 2, as the receiver $Y_{1}$ observes the channel noise only, one might think that a cooperative scheme in which $Y_{1}$ applies quantize-bin-and-forward and $Y_{2}$ applies decode-and-forward is the right strategy.
 However, it turns out that such a scheme is not optimal when the message to $Y_{2}$ is a degraded version of the message to $Y_{1}$.
@@
 As far as we know, the above results are the first cases where a combination of quantize-bin-and-forward and decode-and-forward strategies yields an optimal bidirectional cooperation protocol.
-There is no previously known capacity result in the literature for channels with bidirectional cooperation between the receiver.
+There is no previously known capacity result in the literature for channels with bidirectional cooperation between the receivers.
 Moreover, these results demonstrate that a one-round cooperation scheme is sufficient for achieving capacity for this class of channels.
@@
 Proof:
-The achievability is derived by setting $U=X$, $W=V$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$.
-For the converse, note that due to the Markov chain condition $U\rightarrow X\rightarrow Y_{1},Y_{2}$, the more-capable condition implies that
+The achievability is derived by setting $U=X$, $W=V$, $\hat{Y}_{1}=\varnothing$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$.
+Note that the more capable condition and one-sided conferencing imply that the private message for user 2 can always be decoded by user 1.
+For the converse, note that due to the Markov chain condition $U\rightarrow X\rightarrow Y_{1},Y_{2}$, the more-capable condition implies
 $I(X;Y_{2}|U)\leq I(X;Y_{1}|U)\quad\forall P(u,x).$ (76)
@@
 ∎
-Note that the structure of the novel outer bound $\mathcal{R}_{o}$ given in Theorem 1, in particular, with respect to the constraints (9) and (10), is crucial for deriving the capacity results in Theorems 4 and 5.
-In the next section, we show that the bounds can also be used to derive capacity results for certain classes of Gaussian BCs.
+Remark 7.
+The structure of the novel outer bound $\mathcal{R}_{o}$ given in Theorem 1, in particular, with respect to the constraints (9) and (10), is crucial for deriving the capacity results in Theorems 4 and 5.
+Specific combinations of conditions, such as semi-determinism, degraded message sets, and the more capable channel, make it possible for the converse to be tight.
 IV Gaussian BC with Conferencing Decoders
@@
 $\Psi_{2}$ $=\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab}{1-\lambda^{2}}\right)P\right).$ (90)
-The set $\mathcal{R}_{o}^{G}$ constitutes an outer bound on the capacity region of the BC.
+Then $\mathcal{R}_{o}^{G}$ constitutes an outer bound on the capacity region of the BC.
 Proof:
@@
 $\frac{a^{2}+b^{2}-2\lambda ab}{1-\lambda^{2}}=a^{2}.$ (91)
-Note that in the case of $\lambda=1$ (with $a=b$), or $\lambda=-1$ (with $a=-b$), the above equality is to be interpreted in the limiting sense.
+Note that in the case of $\lambda=1$ (with $a=b$), or $\lambda=-1$ (with a=ba=-b), the above equality is to be interpreted in the limiting sense.
 Further, when $\lambda=\frac{b}{a}$, the BC is physically degraded, for which the outer bound $\mathcal{R}_{o}^{G}$ is achievable.
 (See Remark 12.)
-We now discuss the BC with perfectly correlated noises, i.e., $\lambda=1$ or $-1$, but with the assumption that $\lambda\neq\frac{b}{a}$.
+We now discuss the BC with perfectly correlated noises, i.e., $\lambda=1$ or 1-1, but with the assumption that $\lambda\neq\frac{b}{a}$.
 IV-B Capacity Regions for Specific Channels
@@
 Note that by this choice of random variables, when $|\lambda|=1$, we have
-$I(\hat{Y}_{2};Y_{2}|W,U,Y_{1})=\frac{1}{2}\psi\left(\frac{1-\lambda^{2}}{\hat{\sigma}^{2}}\right)=0$ (95)
+$I(\hat{Y}_{2};Y_{2}|W,U,Y_{1})=\psi\left(\frac{1-\lambda^{2}}{\hat{\sigma}^{2}}\right)=0$ (95)
 Moreover,
-$I(X;Y_{1},\hat{Y}_{2})$ $=\frac{1}{2}\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
-$=\frac{1}{2}\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right)$ (96)
+$I(X;Y_{1},\hat{Y}_{2})$ $=\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
+$=\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right)$ (96)
 and
-$I(X;Y_{1},\hat{Y_{2}}|W)$ $=\frac{1}{2}\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
-$=\frac{1}{2}\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right).$ (97)
+$I(X;Y_{1},\hat{Y_{2}}|W)$ $=\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
+$=\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right).$ (97)
 Note that if $|\lambda|=1$ but $\lambda\neq\frac{b}{a}$, then $a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}\neq 0$.
-Therefore, by letting $\hat{\sigma}^{2}\rightarrow 0$, one can make the above two mutual information terms arbitrarily large, so they are would not be in effect in the characterization of the achievable rate region.
+Therefore, by letting $\hat{\sigma}^{2}\rightarrow 0$, one can make the above two mutual information terms arbitrarily large, so they would not be in effect in the characterization of the achievable rate region.
 For the case of $|a|\geq|b|$, the converse proof is readily given by $\mathcal{R}_{o}^{G}$ in Theorem 6.
@@
 $\begin{cases}Y_{1}=X+Z\\ Y_{2}=X-Z\end{cases}$ (98)
-where $Z$ is a zero mean unit variance Gaussian noise.
+where $Z$ is a zero-mean unit-variance Gaussian noise.
 In this scenario, the two receivers see exactly the same noise but with a different sign.
-For this channel, the capacity result of Theorem 7 reduces to the following:
+The capacity region of this channel with degraded message sets can be obtained from Theorem 7 as:
 $R_{0}$ $\leq\psi(P)+C_{12}$ (99)
@@
 Observe that in the case of common message only, even as $P\rightarrow 0$, it is possible to achieve a strictly positive rate of $R_{0}=\min\{C_{12},C_{21}\}$.
-This makes concrete the statement made in the introduction, namely, that it is possible to use an infinitesimally small amount of power to transmit information at a strictly positive rate.
+This makes concrete the statement made in the introduction of this paper, namely, that it is possible to use an infinitesimally small amount of power to transmit information at a strictly positive rate.
 Note that the argument presented in the proof of Theorem 7 is only valid for strictly positive values of input power $P$.
@@
 Remark 10.
-For the special case of primitive relay channel with perfectly correlated noise, where $Y_{2}$ acts as a relay for $Y_{1}$, $C_{12}=0$ and there is only a private message for the first receiver, the capacity result given in Theorem 7 is reduced to $R_{1}\leq\psi\left(a^{2}P\right)+C_{21}$.
+For the special case of a primitive relay channel with perfectly correlated noises, i.e., $|\lambda|=1$, where $Y_{2}$ acts as a relay for $Y_{1}$, $C_{12}=0$, and there is only a private message for the first receiver, the capacity result given in Theorem 7 is reduced to $R_{1}\leq\psi\left(a^{2}P\right)+C_{21}$.
 In this case, using the quantize-bin-and-forward strategy, one can achieve the cut-set bound.
-Interestingly, the capacity does not depend on $b$ at all.
-This result is an example of a semi-deterministic primitive relay channel, because the relay observation $Y_{2}=bX+Z_{2}$ is a deterministic function of the input $X$ and the receiver observation $Y_{1}=aX+Z_{1}$, when $Z_{1}$ and $Z_{2}$ are perfectly correlated and $\lambda\neq\frac{b}{a}$.
+Interestingly, the capacity does not depend on $b$ at all (as long as $\lambda\neq\frac{b}{a}$).
+This result is an example of a semi-deterministic primitive relay channel, because the relay observation $Y_{2}=bX+Z_{2}$ is a deterministic function of the input $X$ and the receiver observation $Y_{1}=aX+Z_{1}$, when $Z_{1}$ and $Z_{2}$ are perfectly correlated (i.e., $|\lambda|=1$) and $\lambda\neq\frac{b}{a}$.
 In this case, the cut-set bound is achievable [33].
@@
 $\begin{cases}Y_{1}=X+Z\\ Y_{2}=Z\end{cases}$ (103)
-where $Z$ is a zero mean unit variance Gaussian noise.
+where $Z$ is a zero-mean unit-variance Gaussian noise.
 In this case, the receiver $Y_{2}$ does not receive any information from the transmitter and only observes the additive noise of the receiver $Y_{1}$.
-For this channel, the capacity region can be derived by setting $a=1$ and $b=0$ in (92)-(94) and is given by
+The capacity region of this channel with degraded message sets can be derived by setting $a=1$, $b=0$ in (92)-(94) as:
 $R_{0}$ $\leq C_{12}$ (104)
@@
 $\begin{cases}Y_{1}=Z\\ Y_{2}=X+Z.\end{cases}$ (106)
-In this case, the capacity region is given by
+The capacity region of this channel with degraded message sets is given by
 $R_{0}$ $\leq\psi(P)+C_{12}$ (107)
@@
 Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal, which is in fact just the channel noise) then forwards the bin index of the decoded common message to the receiver $Y_{2}$ through the link $C_{12}$.
-Lastly, the receiver $Y_{2}$ decodes the common message using the bin index received and its received signal.
+Lastly, the receiver $Y_{2}$ decodes the common message using the received bin index and its own received signal.
 Note that in Example 5, as the receiver $Y_{1}$ observes the channel noise only, one might think that a cooperative scheme in which $Y_{1}$ applies quantize-bin-and-forward first followed by $Y_{2}$ applying decode-and-forward should be used.
@@
 Theorem 7 is a capacity result for the Gaussian BC with degraded message sets, and with bidirectional conferencing decoders and perfectly correlated noises.
-Next, we establish the capacity region of a Gaussian BC with both common and private messages, and with one-sided cooperation.
+Next, we treat the case of a Gaussian BC with both common and private messages, but with one-sided cooperation.
 Theorem 8.
@@
 for some $\beta\in[0,1]$.
-The proof of Theorem 8 is similar to that of Theorem 7 and therefore omitted.
+Proof:
+The proof is similar to that of Theorem 7. ∎
 For the Gaussian BC in which the noises are not perfectly correlated, the inner and outer bounds of this paper can provide approximate capacity results.
@@
 $R_{0}+R_{1}$ $\leq\psi\left(\beta a^{2}P\right)+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$
-+$\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}+C_{12}$ (114)
+$\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}+C_{12}$ (114)
 $R_{0}+R_{1}$ $\leq\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+(1-\lambda^{2})a^{2}}{2(1-\lambda^{2})}\right)P\right)$
@@
 Proof:
-The above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
+The above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{1}=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
 By a comparison, one can see that the right-hand sides of the constraints (111), (112), (113), (114) and (115) are within half a bit of (83), (81), (88), (85), and (86), respectively. ∎
-Next, we present an approximate capacity result for the Gaussian BC with one-way conferencing.
+Next, we present an approximate capacity result for the Gaussian BC with one-sided conferencing link.
 Theorem 10.
@@
 $R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\beta a^{2}P\right)+\psi\left(\frac{(1-\beta)b^{2}P}{\beta b^{2}P+1}\right)$
-+$\qquad\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}$ (118)
+$\qquad\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}$ (118)
 $R_{0}+R_{1}+R_{2}$ $\leq\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+(1-\lambda^{2})a^{2}}{2(1-\lambda^{2})}\right)P\right)$
@@
 Proof:
-Similar to Theorem 9, the above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
+Similar to Theorem 9, the above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{1}=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
 By a simple comparison, one can see that the right-hand sides of the constraints (116), (117), (118) and (119) are within half a bit of (83), (88), (85), and (86), respectively. ∎
@@
 Proof:
-The bound $\mathcal{R}_{i}^{DF-G}$ is derived from $\mathcal{R}_{i}^{(1)}$ given in Theorem 2 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{2}=\varnothing$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $\beta P$ and $(1-\beta)P$, respectively.
+The bound $\mathcal{R}_{i}^{DF-G}$ is derived from $\mathcal{R}_{i}^{(1)}$ given in Theorem 2 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{1}=\hat{Y}_{2}=\varnothing$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $\beta P$ and $(1-\beta)P$, respectively.
 Note that this inner bound is in fact derived for the channel with one-sided cooperation (i.e., it does not make use of the conferencing link $C_{21}$) using the decode-and-forward strategy alone.
@@
 Thus, $\mathcal{R}_{i}^{DF-G}$ is the capacity region.
-As a concluding remark for this section, the novel structure of the outer bound given in Theorem 6, in particular, the constraint (86), is crucial for deriving the exact capacity results in Theorems 7 and 8, and also the approximate capacity results in Theorems 9, 10, and 11.
+As a concluding remark for this section, the novel structure of the outer bound given in Theorem 6, in particular, the constraint (86), is crucial for deriving the exact capacity results in Theorems 7 and 8, and also the approximate capacity results in Theorems 9, 10, and $1$.
 V Conclusions
@@
 Appendix A Proof of Theorem 1
-We only derive the constraints involving the auxiliary variable $U$, i.e., (1), (5), (6), (8), and (10).
-The constraints involving $V$ can be derived symmetrically.
-The last constraint is due to the cut-set bound.
+We only derive the constraints that include the auxiliary variable $U$, i.e., (1), (5), (6), (8), and (10).
+The constraints that include $V$ can be derived symmetrically.
+The last constraint (11) is due to the cut-set bound.
 Consider a length-$n$ code with common and private message rates $(R_{0},R_{1},R_{0})$ and with vanishing average probability of decoding error. Define the auxiliary variables as follows:
@@
 By substituting (137) and (139) into (132), we obtain the desired bounds on $R_{2}$ (5) and (6).
-Next, we derive the constraints on the sum rate:
+Next, we derive the constraints on the sum rate, starting from Fano’s inequality:
 $n$ $(R_{0}+R_{1}+R_{2})$
@@
 […]
-Lastly, (11) is just the cut-set bound. By applying a standard time-sharing argument, the proof is thus complete. The cardinality bounds on $U$ and $V$ are due to Carathéodory theorem.
+By applying a standard time-sharing argument, the proof is thus complete. The cardinality bounds on $U$ and $V$ are due to Carathéodory theorem [29].
 Appendix B Proof of Proposition 1
@@
 and let $\alpha_{1}\in[0,1]$.
-We construct random codebooks consisting of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively.
+We construct random codebooks consisting of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $1:2^{nR_{0}}]$, $1:2^{nR_{1}}]$, and $1:2^{nR_{2}}]$, respectively.
 First of all, each of the private messages $M_{1}$ and $M_{2}$ is split into two parts as follows:
@@
     $(W^{n}(m_{00}),U^{n}(m_{00},m_{11},b_{1}^{\mathcal{T}}),V^{n}(m_{00},m_{22},b_{2}^{\mathcal{T}}))\\ \qquad\in\mathcal{T}_{\epsilon}^{n}(P_{WUV}).$ (172)
-    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (172), let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}+b_{2}$.
-If there is no such a pair, declare an encoding error.
+    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (172), let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}$.
+If multiple such pairs have the same minimum $b_{1}$, choose the one with the minimum $b_{2}$.
+If no pair $(b_{1},b_{2})$ satisfies (172), declare an encoding error.
 Based on the mutual covering lemma [29], one can easily see that for sufficiently large $n$ and small $\epsilon$, the probability of encoding error tends to zero provided that
@@
   2. 2.
-    Partition the set $[1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21})}$ elements.
+    Partition the set $1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21})}$ elements.
 For a given $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$, let $\mathfrak{C}_{21}(\hat{m}_{2})$ denote the index of the bin that $\hat{m}_{2}$ belongs to.
   3. 3.
-    Given the output sequence $Y_{2}^{n}$ at the second receiver, let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies
+    Given the output sequence $Y_{2}^{n}$ at the second receiver, let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $1:2^{n\hat{R}_{2}}]$ that satisfies
     $(Y_{2}^{n},\hat{Y}_{2}^{n}(\hat{m}_{2}^{\mathcal{T}}))\in\mathcal{T}_{\epsilon}^{n}(P_{Y_{2}\hat{Y}_{2}}).$ (174)
-    If there is no integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies (174), then declare an error. The probability of this error tends to zero provided that
+    If there is no integer in $1:2^{n\hat{R}_{2}}]$ that satisfies (174), then declare an error. The probability of this error tends to zero provided that
     $\hat{R}_{2}\geq I(Y_{2};\hat{Y}_{2}).$ (175)
@@
   6. 6.
-    Partition the set $[1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}^{q}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12}^{q})}$ elements.
+    Partition the set $1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}^{q}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12}^{q})}$ elements.
 For a given $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$, let $\mathfrak{C}_{12}^{q}(\hat{m}_{1})$ denote the index of the bin that $\hat{m}_{1}$ belongs to.
   7. 7.
-    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies
+    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $1:2^{n\hat{R}_{1}}]$ that satisfies
     $(W^{n}(m_{00}^{*}),U^{n}(m_{00}^{*},m_{11}^{*},b_{1}^{*}),Y_{1}^{n},\hat{Y}_{1}^{n}(\hat{m}_{1}^{\mathcal{T}}))\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WUY_{1}\hat{Y}_{1}}).$ (182)
     Note that $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ is the unique triple decoded in Step 4 according to (176)-(177).
-If there is no integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies (182), then declare an error.
+If there is no integer in $1:2^{n\hat{R}_{1}}]$ that satisfies (182), then declare an error.
 The probability of this error event tends to zero provided that
@@
   8. 8.
-    Partition the set $[1:2^{nR_{00}}]$ into $2^{n\hat{C}_{12}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{12}^{d})}$ elements.
+    Partition the set $1:2^{nR_{00}}]$ into $2^{n\hat{C}_{12}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{12}^{d})}$ elements.
 For a given $m_{00}\in[1:2^{nR_{00}}]$, let $\mathfrak{C}_{12}^{d}(m_{00})$ denote the index of the bin that $m_{00}$ belongs to.
@@
 $I(W,U,Y_{1};\hat{Y}_{2})+C_{21}-I(Y_{2};\hat{Y}_{2})=\\ C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}).$ (195)
-One can easily see that if $C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$.
-Otherwise, we keep $C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$ as is.
+One can easily see that if $C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$. Otherwise, we keep $C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$ as is.
 Then, (B-C)-(194) are equivalent to
@@
 $R_{00}+R_{22}$ $+B_{2}\leq\min\Big{\{}I(W,V;Y_{2})$
-+$+\{\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}+\bar{\alpha}_{1}{C}_{12},$
+$+\{\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}+\bar{\alpha}_{1}{C}_{12},$
 $\ I(W,V;\hat{Y}_{1},Y_{2})+\bar{\alpha}_{1}{C}_{12}\Big{\}}.$ (204)
-Lastly, by considering the bounds (173), (B-C)-(197), and (B-C)-(204) and by applying the Fourier–Motzkin algorithm to eliminate all the partial rates, one can derive the rate region $\mathcal{R}_{i}^{(1)}$.
+Lastly, considering the bounds in (173), (B-C)-(197), and (B-C)-(204) together, by applying the Fourier–Motzkin algorithm to eliminate all the partial rates, one can derive the rate region $\mathcal{R}_{i}^{(1)}$.
 The proof is thus complete.
-We note that the error analysis here can also be derived from the general inner bounds of [33], but the general expressions of [33] are not always easy to simplify.
+We note that the error probability analysis here can also be derived from the general approach of [34], but the general expressions in [34] are not always easy to simplify.
 Appendix C Proof of Proposition 2
@@
 and let $\alpha_{2}\in[0,1]$.
-Similar to the proof of Proposition 1, we construct random codebooks of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively.
+Similar to the proof of Proposition 1, we construct random codebooks of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $1:2^{nR_{0}}]$, $1:2^{nR_{1}}]$, and $1:2^{nR_{2}}]$, respectively.
 Each of the private messages $M_{1}$ and $M_{2}$ is split into two parts as follows:
@@
 $R_{2}=R_{20}+R_{22}.$ (208)
-The sub-messages $M_{10}$ and $M_{20}$ are decoded at both receivers, i.e., the triple $(M_{0},M_{10},M_{20})$ is considered as a common message with rate
+The sub-messages $M_{10}$ and $M_{20}$ are decoded at both receivers, i.e., the triple $(M_{0},M_{10},M_{20})$ is considered as the common message with rate
 $R_{00}=R_{0}+R_{10}+R_{20}.$ (209)
@@
     $(W^{n}(m_{00}),U^{n}(m_{00},m_{11},b_{1}^{\mathcal{T}}),V^{n}(m_{00},m_{22},b_{2}^{\mathcal{T}}))\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WUV}).$ (213)
-    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (213), then let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}+b_{2}$.
-If there is no such a pair, declare encoding error.
+    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (213), let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}$.
+If multiple such pairs have the same minimum $b_{1}$, choose the one with the minimum $b_{2}$.
+If no pair $(b_{1},b_{2})$ satisfies (213), declare an encoding error.
 Based on the mutual covering lemma [29], one can easily see that for sufficiently large $n$ and small $\epsilon$, the probability of encoding error tends to zero provided that
@@
   1. 1.
-    Partition the set $[1:2^{nR_{00}}]$ into $2^{n\hat{C}_{21}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{21}^{d})}$ elements.
+    Partition the set $1:2^{nR_{00}}]$ into $2^{n\hat{C}_{21}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{21}^{d})}$ elements.
 For a given $m_{00}\in[1:2^{nR_{00}}]$, let $\mathfrak{C}_{21}^{d}(m_{00})$ denote the index of the bin that $m_{00}$ belongs to.
@@
   3. 3.
-    Partition the set $[1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}^{q}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21}^{q})}$ elements.
+    Partition the set $1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}^{q}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21}^{q})}$ elements.
 For a given $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$, let $\mathfrak{C}_{21}^{q}(\hat{m}_{2})$ denote the index of the bin that $\hat{m}_{2}$ belongs to.
@@
   5. 5.
-    Let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies
+    Let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $1:2^{n\hat{R}_{2}}]$ that satisfies
     $(W^{n}(m_{00}^{\bullet}),Y_{2}^{n},\hat{Y}_{2}^{n}(\hat{m}_{2}^{\mathcal{T}}))\in\mathcal{T}_{\epsilon}^{n}(P_{WY_{2}\hat{Y}_{2}}).$ (217)
-    If there is no integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies (217), then declare an error. The probability of this error tends to zero provided that
+    If there is no integer in $1:2^{n\hat{R}_{2}}]$ that satisfies (217), then declare an error. The probability of this error tends to zero provided that
     $\hat{R}_{2}\geq I(W,Y_{2};\hat{Y}_{2})$ (218)
@@
   8. 8.
-    Partition the set $[1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12})}$ elements.
+    Partition the set $1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12})}$ elements.
 For a given $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$, let $\mathfrak{C}_{12}(\hat{m}_{1})$ denote the index of the bin that $\hat{m}_{1}$ belongs to.
   9. 9.
-    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies
+    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $1:2^{n\hat{R}_{1}}]$ that satisfies
     $(W^{n}(m_{00}^{*}),U^{n}(m_{00}^{*},m_{11}^{*},b_{1}^{*}),Y_{1}^{n},\hat{Y}_{1}^{n}(\hat{m}_{1}^{\mathcal{T}}))\\ \in\mathcal{T}_{\epsilon}^{n}(P_{WUY_{1}\hat{Y}_{1}}),$ (226)
     where $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ is the unique triple decoded in Step 6 according to (219)-(221).
-If there is no integer $[1:2^{n\hat{R}_{1}}]$ that satisfies (226), then declare an error.
+If there is no integer $1:2^{n\hat{R}_{1}}]$ that satisfies (226), then declare an error.
 The probability of this error event tends to zero provided that
@@
 $I(W,U,Y_{1};\hat{Y}_{2})+{\alpha}_{2}C_{21}-I(W,Y_{2};\hat{Y}_{2})=\\ {\alpha}_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1}).$ (236)
-One can easily see that if ${\alpha}_{2}C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$. Thus, (232)-(235) are equivalent to
+One can easily see that if ${\alpha}_{2}C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$.
+Thus, (232)-(235) are equivalent to
 $R_{11}+B_{1}$ $\leq\min\Big{\{}I(U;Y_{1}|W)$
-+$\qquad+\{{\alpha}_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+},$
+$\qquad+\{{\alpha}_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+},$
 $\qquad\qquad I(U;Y_{1},\hat{Y}_{2}|W)\Big{\}}$ (237)
@@
 $R_{22}+B_{2}$ $\leq\min\Big{\{}I(V;Y_{2}|W)$
-+$\qquad+\{C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+},$
+$\qquad+\{C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+},$
 $\quad\qquad I(V;\hat{Y}_{1},Y_{2}|W)\Big{\}}$ (243)
 $R_{00}+R_{22}+B_{2}$ $\leq I(W,V;Y_{2}).$ (244)
-Finally, by considering the bounds (237)-(238) and (243)-(244) and by applying the Fourier–Motzkin elimination algorithm, we obtain the rate region $\mathcal{R}_{i}^{(2)}$. The proof is thus complete.
+Finally, considering the bounds (237)-(238) and (243)-(244) together, by applying the Fourier–Motzkin algorithm to eliminate all the partial rates, we obtain the rate region $\mathcal{R}_{i}^{(2)}$.
+The proof is thus complete.
 Appendix D Proof of Theorem 6
@@
 First note that since $|a|\geq|b|$, the receiver $Y_{1}$ is less noisy than the receiver $Y_{2}$, i.e.,
-$I(V;Y_{2})\leq I(V;Y_{1}).$ (245)
+$I(V;Y_{2})\leq I(V;Y_{1})\qquad\forall P(u,x).$ (245)
 Therefore, we have
@@
 $I_{4}$ $\doteq I(X;Y_{1},Y_{2}).$ (254)
-Consider first the special cases where $\lambda=\frac{b}{a}$. In this case, $Z_{1}$ and $Z_{2}$ can be thought of as being related as
+Consider first the special case where $\lambda=\frac{b}{a}$. In this case, $Z_{1}$ and $Z_{2}$ can be thought of as being related as
 $Z_{2}=\frac{b}{a}Z_{1}+\tilde{Z}$, (255)
-where $Z_{1}$, $\tilde{Z}$ are independent. Since $Y_{1}=aX+Z_{1}$, we have
+where $_{1}, Z~\tilde{Z}$ are independent. Since $Y_{1}=aX+Z_{1}$, we have
 $Y_{2}$ $=bX+Z_{2}$ (256)
@@
 $=\frac{b}{a}Y_{1}+\tilde{Z}$. (258)
-So, the BC is physically degraded, i.e., $X\rightarrow Y_{1}\rightarrow Y_{2}$ forms a Markov chain.
-The simultaneous maximization of the four mutual information terms (251)-(254) then reduces to the simultaneous maximization of $h(Y_{1}|V)$ and $h(Y_{1})$ together with minimization of $h(Y_{2}|V)$, which is achieved by jointly Gaussian $(X,V)$ as can be shown by the entropy power inequality for the degraded Gaussian broadcast channel [29].
+So, the BC is physically degraded (because $|a|\geq|b|$), i.e., $X\rightarrow Y_{1}\rightarrow Y_{2}$ forms a Markov chain.
+In this special case, the simultaneous maximization of the four mutual information terms (251)-(254) then reduces to the simultaneous maximization of $h(Y_{1}|V)$ and $h(Y_{1})$ together with minimization of $h(Y_{2}|V)$, which is achieved by jointly Gaussian $(X,V)$ as can be shown by the entropy power inequality in the proof of the capacity region of the degraded Gaussian broadcast channel [29].
 In the rest of the proof, we assume $\lambda\neq\frac{b}{a}$.
@@
 $I(X;Y_{1},Y_{2}|V)$ $=I(X;\hat{Y},Y_{2}|V)=I(X;\hat{Y}|V),$ (265)
-$I(X;Y_{1},Y_{2})$ $=I(X;\hat{Y},Y_{2})=I(X;\hat{Y}).$ (266)
-Now, utilizing (265)-(266) for bounding the maximum values of (251)-(254), we see that after expressing the mutual information in terms of differential entropy, the optimization is effectively to maximize $h(\hat{Y}|V)$, $h(Y_{1}|V)$ and $h(\hat{Y})$, and simultaneously to minimize $h(Y_{2}|V)$.
+[…]
+Now, utilizing (265)-(266) for bounding the maximum values of (251)-(254), we see that after expressing the mutual information in terms of differential entropy, the optimization is effectively to maximize $h(\hat{Y}|V)$, $h(Y_{1}|V)$ and h(Y^)h($\hat${Y}), and simultaneously to minimize $h(Y_{2}|V)$.
 We complete this step using the entropy power inequality. First,
@@
 We then have
-$h(\hat{Y}|V)$ $=h(\mu X+\hat{Z})$ (273)
+[…]
 $=h\left(aX+\frac{a}{\mu}\hat{Z}\right)-\frac{1}{2}\log\left(\frac{a^{2}}{\mu^{2}}\right)$ (274)
@@
 where (281) is due to the entropy power inequality, and (282) holds because $Z_{2}$ has the same distribution as $Z_{1}$ and both are independent of $(X,V)$, and also $Z_{2}^{*}$ is independent of $V$.
-Considering (272), (278), and (284) together with the fact that a Gaussian $X\sim\mathcal{N}(0,P)$ maximizes $h(\hat{Y})$ allows us to conclude that a jointly Gaussian $(X,V)$ simultaneously maximizes the mutual information terms in (251)-(254) and (265)-(266).
+Considering (272), (278), and (284) together with the fact that a Gaussian $X\sim\mathcal{N}(0,P)$ maximizes h(Y^)h($\hat${Y}) allows us to conclude that a jointly Gaussian $(X,V)$ simultaneously maximizes the mutual information terms in (251)-(254) and (265)-(266).
 The proof can be completed by substituting the evaluation of these mutual information terms (251)-(254) for Gaussian $(X,V)$ into $\mathcal{R}_{o}$ in Theorem 1.